### v1

# MIST: Mutual Information Estimation via Supervised Training

Abstract
We propose a fully data-driven approach to designing mutual information (MI) estimators. Since any MI estimator is a function of the observed sample from two random variables, we parameterize this function with a neural network (MIST) and train it end-to-end to predict MI values. Training is performed on a large meta-dataset of 625,000 synthetic joint distributions with known ground-truth MI. To handle variable sample sizes and dimensions, we employ a two-dimensional attention scheme ensuring permutation invariance across input samples. To quantify uncertainty, we optimize a quantile regression loss, enabling the estimator to approximate the sampling distribution of MI rather than return a single point estimate. This research program departs from prior work by taking a fully empirical route, trading universal theoretical guarantees for flexibility and efficiency. Empirically, the learned estimators largely outperform classical baselines across sample sizes and dimensions, including on joint distributions unseen during training. The resulting quantile-based intervals are well-calibrated and more reliable than bootstrap-based confidence intervals, while inference is orders of magnitude faster than existing neural baselines. Beyond immediate empirical gains, this framework yields trainable, fully differentiable estimators that can be embedded into larger learning pipelines. Moreover, exploiting MI’s invariance to invertible transformations, meta-datasets can be adapted to arbitrary data modalities via normalizing flows, enabling flexible training for diverse target meta-distributions.
1 Introduction
Mutual information (MI) is a measure of nonlinear statistical dependence between two random variables, quantifying how much knowing one variable reduces the uncertainty about the other (shannon1948mathematical; mackay2003information). For two continuous random variables $X$ and $Y$, the MI is written in terms of their joint distribution $P_{(X,Y)}$ and marginal distributions $P_{X}$ and $P_{Y}$:
$\operatorname{I}(X;Y)=\int_{\mathcal{Y}}\int_{\mathcal{X}}P_{(X,Y)}(x,y)\log\left(\frac{P_{(X,Y)}(x,y)}{P_{X}(x)\,P_{Y}(y)}\right)\,dx\,dy.$
Mutual information is one of the most widely applied information measures in data science, finding applications in feature selection (peng2005feature_selection; kwak2002_feature_selection), causality (butte1999_causality_genomics), representation learning (chen2016infogan; higgins2018towards_distentagled_representations; tishby2015_info_bottleneck; zhao2018information_autoencoders), reinforcement learning (oord2018_cpc; pathak2017curiosity), and generative modeling (alemi2016deep_variation_info_bottleneck; alemi2018gilbo; huang2020evaluating_lossy_compression).
Calculating MI requires access to the joint and marginal distributions, which are rarely known in practice. Consequently, most approaches infer MI from finite samples drawn from these distributions. Existing estimators fall into two main families. (1) Density estimators approximate the underlying densities, for example, using kernel density estimation, $k$-nearest neighbors, or variational autoencoders. MI can then be computed by substituting these estimates into one of MI’s definitions (such as Equation 1). Prominent density estimators include KSG (ksg) and MINDE (minde). (2) Density relationship estimators instead estimate the relationship between the joint and marginal distributions, often using variational lower bounds derived from the Donsker–Varadhan representation of the Kullback–Leibler divergence (see Appendix A.1 for these definitions). Notable examples include MINE (mine), NWJ (nwj), and SMILE (MINE_exp_variance_consistency_tests).
In both families of approaches, neural methods have been developed to improve the estimator’s capacity to capture complex density functions or density relationships. Broadly, these neural methods have achieved better performance on moderately complex distributions and high-MI settings. However, both families of approaches struggle in challenging settings that are common in application, including high-dimension, limited samples, complex distributions, and high MI (bmi_benchmark; mcallester2020formal_limitations). Prominent examples of these settings include applications in genomics (klein2015droplet_genomics), neuroscience (ganguli2012compressed_neuroscience), astronomy (pandey2017much_astronomy), and machine learning (goldfeld2018estimating_info_flow). While previous benchmarking efforts have expanded the set of distribution families considered, MI estimators have rarely been empirically evaluated under these challenging conditions individually, and even more rarely in combination. Most empirical evaluations have studied estimation for MI below 1, for fewer than 5 distributions, with a minimum sample size above 10,000, and for higher dimensions limited to normal distributions. See Appendix A.3 for an overview of the empirical evaluation scale for existing methods.
In this work, we propose a fully data-driven, empirical approach to designing mutual information estimators. Instead of building estimators that compute MI by approximating density functions or their relation, we learn to predict MI directly from data. We parameterize a mutual information estimator as a neural network and train such a network end-to-end on a large meta-dataset of synthetic joint distributions with known ground-truth MI values. Our proposed model, MIST (Mutual Information estimation via Supervised Training), handles arbitrary sample sizes and dimensions, and features built-in uncertainty estimates. This meta-learning approach takes inspiration from amortized inference procedures such as simulation-based inference (SBI; pmlr-v89-papamakarios19a; cranmer2020frontier), amortized Bayesian inference (ABI; gonccalves2020training; elsemueller2024sensitivity; radev2023bayesflow; avecilla2022neural; pmlr-v235-gloeckler24a), and meta-statistical learning (peyrard2025meta).
Our work focuses on the setting for MI estimators that was underexplored in previous work. Empirically, we conduct a large-scale evaluation that systematically explores the performance of our estimator in the most difficult regimes: with fewer samples (10-500), in higher dimensions (2-32), and across a wider range of MI values (0-40). We find that our learned estimator outperforms existing methods across these conditions, especially in low-sample regimes. We provide an empirical study of the behavior of estimators under increasing dimensionality, finding that our models require roughly half as many samples on average to provide reliable MI estimates as the best baseline. The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives, while inference is orders of magnitude faster than existing neural methods. Finally, the learned estimators generalize to unseen distributions and sample sizes.
Beyond empirical performance, this meta-statistical framework establishes a new paradigm for constructing information-theoretic estimators. Because the estimator is fully differentiable and trainable, it can be embedded in larger learning pipelines, fine-tuned to specific data modalities, and systematically improved through curated meta-datasets. Leveraging MI’s invariance under invertible transformations, synthetic training distributions can, in the future, be adapted to arbitrary domains and modalities via normalizing flows (papamakarios2021normalizing_flows), paving the way for general-purpose, learned estimators of mutual information.
We provide an open-source library [^1] https://github.com/grgera/mist for training and evaluating meta-learned MI estimators, and include a detailed discussion of these opportunities and future work in Section 5.
2 Related Work
Mutual information estimation is a foundational challenge across data science tasks, for which a wide variety of estimation methods have been developed. Although these approaches vary in their mathematical foundations and computational strategies, they can be grouped into two main families: density estimators and density ratio estimators.
Density Estimators.
Density estimators compute mutual information estimates by performing intermediate estimates of density functions. The most straightforward approach is to use MI’s analytical definition, which requires estimates of the joint and marginal density functions. Histogram MI estimators approximate these density functions by partitioning the support of each variable into a set of bins and counting the frequencies in each (histogram_binning). Histogram estimators suffer from very poor scaling and are highly sensitive to the choice of bins, leading to research on optimizing the bin selection (adaptive_binning). Kernel density estimators (KDE; kde; textbook_kde) were later developed to help avoid this sensitivity by replacing fixed bins with kernel functions centered at each sample point.
original_kl_density_estimator designed a method for entropy estimation using neighborhood distances, an approach which was later refined with weighted variants (weighted_kl_density_estimator). The Kraskov-Stögbauer-Grassberger estimator (KSG; ksg) extended this k-NN framework to mutual information estimation using MI’s expression in terms of conditional entropy (see Appendix A.1). KSG is one of the most broadly used MI estimators, especially for low-sample settings, and serves as a primary benchmark comparison in this work.
Several approaches have explored the use of deep generative models for density estimation. One line of work uses normalizing flows to reconstruct density functions directly. However, these methods suffer from poor estimates even for data with simple structure (MINE_exp_variance_consistency_tests; dine_gaussian). The Barber-Agakov bound (BA; barber2004algorithm_BA_Bound) provides a variational lower bound on MI by introducing a variational approximation $q(x|y)$ to the true conditional $p(x|y)$. Several works have applied this bound using conditional generative models with different choices for the marginal distribution (alemi2016deep_variation_info_bottleneck; chalk2016relevant_BA_bound_application; kolchinsky2019nonlinear). Later work used VAEs to learn the joint and marginal density functions directly (MINE_exp_variance_consistency_tests). Most recently, MINDE (minde) trains conditional or joint diffusion models to learn score functions, then applies the Girsanov theorem to compute KL-divergence as the expected difference between joint and marginal scores. VCE (chen2025neural_vector_copulas) is a concurrent work to ours, and writes MI as the negative differential entropy of the vector copula density. Their approach uses flow models to estimate the marginal densities, then estimates the vector copula density using maximum likelihood estimation.
Another line of density estimators has focused on building MI estimators for complex settings by learning methods to simplify the distributions before applying classical estimators or computing MI directly. MIENF (mi_through_normalizing_flows) learns a pair of normalizing flows that apply MI-preserving transformations to the joint distribution, enabling closed-form MI calculation. Latent-MI (LMI; lmi_estimator) trains a network to compress data with minimal mutual information loss, then applies KSG to the low-dimensional embeddings. These methods have improved abilities to handle high-dimensional data, but have large sample and computational requirements, and rely on the existence of simple or low-dimensional structure in high-dimensional data.
Density Ratio Estimators.
Density ratio estimators compute mutual information by estimating the relationship between the joint density $P(X,Y)$ and the product of marginal densities $P_{X}P_{Y}$. Since mutual information can be equivalently expressed as the Kullback-Leibler divergence $D_{KL}(P(X,Y)\|P_{X}P_{Y})$ (donsker_varadhan_kl_div_mi_definition), many density ratio methods leverage variational bounds on KL divergence. Several methods have been developed by defining discriminative tasks that implicitly predict this density ratio. MINE (mine) trains a neural network critic to differentiate joint and marginal samples, optimizing the critic to maximize the Donsker-Varadhan (DV) lower bound. The Nguyen-Wainwright-Jordan estimator (NWJ; nwj) trains a critic model similarly, but uses an alternative variational bound which offers lower variance at the cost of a looser bound. InfoNCE (oord2018_cpc) trains a classifier critic to identify true joint samples among negative samples drawn from the marginals. SMILE (MINE_exp_variance_consistency_tests) combines the InfoNCE and NWJ bounds through a clip operation, interpolating between their respective bias-variance profiles to achieve more flexible trade-offs. The limitations of variational estimators are well-studied in mcallester2020formal_limitations, MINE_exp_variance_consistency_tests, and poole_variation_bounds_survey, and include high variance estimates, high sample requirements, and poor estimates for high MI settings. We refer the reader to poole_variation_bounds_survey for a detailed review of MI variational bounds.
Evaluation of MI Estimators.
Evaluating MI estimators is itself challenging: it requires distributions with known ground-truth MI and large sample sizes (required by most methods). Existing work has therefore primarily focused on theoretical guarantees of low bias, consistency with increasing samples, and the tightness of variational bounds (mine; nwj; MINE_exp_variance_consistency_tests). Empirical evaluations have been limited in scope. Early methods, such as binning and kernel density estimators, were tested on simple time-series data like sine waves (kde). Later neural methods began evaluation on a smaller set of synthetic distributions: MINE and SMILE were evaluated on 3–4 distributions, including high-dimensional Gaussians up to 20D (mine; MINE_exp_variance_consistency_tests). The majority of methods have been evaluated exclusively on MI values below 1, often on Gaussian or near-Gaussian distributions (see Table 3). Methods that consider higher dimensionalities do so largely by scaling low-dimensional distributions (assuming low-dimensional latent structure), or test only on high-dimensional Gaussians (lmi_estimator; minde).
Recently, the BMI benchmark (bmi_benchmark) applied 20 invertible transformations to normal and Student’s t-distributions to create a broader set of distribution families with unique challenges, such as sparsity and long tails. This evaluation on more complex distributions highlighted the limitations of many existing estimators, leading authors to argue that multivariate normal distributions provided a overly optimistic view of estimator performance. BMI has been used by some of the most recent estimators, such as MINDE. We provide a description of our datasets in Appendix B, and an overview of existing evaluation scopes in Appendix A.3.
Machine Learning for Statistical Inference.
Our work aligns with the broader research direction on learning inference procedures. A relevant related direction concerns neural processes (garnelo2018neural; garnelo2018conditional; kim2019attentive; Gordon:2020; markou2022practical; huang2023practical; BruinsmaMRFAVBH23), which can predict latent variables of interest from datasets (chang2025amortized) by leveraging transformers (pmlr-v162-nguyen22b) and Deep Sets (NIPS2017_f22e4747) to enforce permutation invariance (JMLR:v21:19-322). A related approach, known as prior-fitted networks, has demonstrated that transformers can be effectively repurposed for Bayesian inference (muller2022transformers) and optimization tasks (pmlr-v202-muller23a). Additionally, there is growing interest in using trained models to assist in statistical inference tasks (angelopoulos2023predictionpoweredinference) and optimization (NIPS2017_addfa9b7; NEURIPS2020_f52db9f7; NEURIPS2021_56c3b2c6; amos2023tutorialamortizedoptimization) like simulation-based inference (SBI; pmlr-v89-papamakarios19a; cranmer2020frontier), amortized Bayesian inference (ABI; gonccalves2020training; elsemueller2024sensitivity; radev2023bayesflow; avecilla2022neural; pmlr-v235-gloeckler24a) and the learning of frequentist statistical estimators (peyrard2025meta).
3 Learning Framework
Figure 1: We propose a fully data-driven, empirical approach to designing mutual information estimators. We design a large empirical meta-dataset composed of samples from a set of distributions with known MI (left), and train a SetTransformer-based model to predict MI directly from sets of samples (right).
The methods available today decompose MI estimation into intermediate tasks of density function or density ratio estimation. To estimate a density function (or ratio) beyond very simple settings, these methods require large amounts of samples as part of inference. Even with access to a high number of samples, density estimation itself is a notoriously difficult and unsolved problem, particularly for high-dimensional or complex distributions (density_estimation). As a result, existing MI estimators still fail in many realistic data settings. They scale poorly to high-dimensional data and complex distributions, are unstable for low-sample settings, and require expensive inference procedures. An additional consequence is that studying and improving MI estimators is challenging, with new methods often resulting from theoretical discoveries.
Our work offers an alternative philosophy for designing mutual information estimators. Instead of building estimators that compute MI with a pre-specified algorithm (e.g., approximating density functions or their ratios), we propose to predict MI directly from data. We take a fully empirical perspective, asking: Can a model learn a generalizable algorithm for predicting mutual information directly from samples?
To learn this algorithm, we parameterize a mutual information estimator as a neural network that we call MIST (Mutual Information estimation from Supervised Training). MIST takes as input a set of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $P_{XY}$, and outputs an estimate of $I(X;Y)$. The model is trained end-to-end on a large meta-dataset of synthetic joint distributions with known ground-truth MI values. We have reflected our framework in Figure 1 and described the meta-dataset design, architecture, and training choices below.
The Learning Task
Let $\Gamma$ denote a family of joint data-generating distributions over $\mathcal{X}\times\mathcal{Y}$. For any $\gamma\in\Gamma$, let $I_{\gamma}$ denote the mutual information between $X$ and $Y$ under $\gamma$. A mutual information estimator observes a finite dataset $D:=\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn i.i.d. from $\gamma$, and aims to estimate $I_{\gamma}$ from samples alone. We parametrize the estimator as a function $f_{\theta}:\bigcup_{n=1}^{\infty}(\mathcal{X}\times\mathcal{Y})^{n}\to\mathbb{R}_{\geq 0},$ where $\theta$ are neural network weights. To train $f_{\theta}$, we assume access to a meta-distribution $P_{\Gamma}$ over generative laws. First, a distribution $\gamma\sim P_{\Gamma}$ is sampled, then a dataset $D$ is drawn from $\gamma$, together with its ground-truth mutual information $I_{\gamma}$. The training objective is to minimize the mean squared error:
$\mathcal{L}_{\mathrm{MSE}}(\theta)=\mathbb{E}\!\left[\,\big(f_{\theta}(D)-I_{\gamma}\big)^{2}\,\right],$ (1)
where the expectation is taken over $D\sim\gamma$ and $\gamma\sim P_{\Gamma}$. In the limit of sufficient model capacity and perfect optimization, $f_{\theta}$ converges to the Bayes-optimal regression function: $f^{*}(D)=\mathbb{E}\!\left[\,I_{\gamma}\,\middle|\,D\,\right].$ That is, the optimal estimator outputs the posterior expectation of the mutual information conditioned on the observed dataset. This shows the soundness of the approach, which is further clarified by the following decomposition of the MSE loss:
$\mathcal{L}_{\mathrm{MSE}}(\theta)=\mathbb{E}\!\left[\big(f_{\theta}(D)-f^{*}(D)\big)^{2}\right]+\mathbb{E}\!\left[\mathrm{Var}\!\left(I_{\gamma}\,\middle|\,D\right)\right],$ (2)
The second term corresponds to the irreducible epistemic uncertainty due to the fact that, for a finite $n$, the dataset $D$ does not uniquely identify the generative law $\gamma$. The irreducible error in MSE comes from the difficulty of the statistical inference task.
Built-in Uncertainty Quantification
[…]
$L_{\tau}(q,\hat{q})=\max(\tau(q-\hat{q}),(1-\tau)(\hat{q}-q)).$ (3)
By querying multiple values of $\tau$, we can approximate the full sampling distribution of MI. Furthermore, these estimates are obtained as the result of a forward pass and are therefore cheap to compute. We evaluate the calibration and reliability of these quantile-based intervals against bootstrap alternatives in Section 4.3 and include both model variants throughout our experiments.
Meta-Dataset
Training our meta-learned estimators in a supervised manner requires distributions with known MI. We generate these distributions using BMI, a library designed to create complex benchmarks for MI estimation. BMI exploits the invariance of mutual information to homeomorphisms, applying invertible transformations to simple base distributions to produce more complex and challenging distributions. We partition the base distribution families into separate train and test sets, that is, the meta-distribution used for training has a different support than the one used for testing. For each base distribution, we apply BMI’s suite of invertible transformations and generate variants across dimensions ranging from $2$ to $32$, yielding two distribution sets $\Gamma_{\text{train}}$ and $\Gamma_{\text{test}}$. From these distributions, we construct a training meta-dataset $\mathcal{M}_{\text{train}}$ and two test meta-datasets differing only in scale: $\mathcal{M}_{\text{test}}$, and $\mathcal{M}_{\text{test-extended}}$. These test meta-datasets contain a mix of distribution families found in $\Gamma_{\text{train}}$ (subsequently referred to as in-meta-distribution or IMD) or unseen at training time (out-of-meta-distribution or OoMD). Since MI is a function of two random variables, each meta-datapoint in our meta-datasets consists of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $\gamma\in\Gamma$, along with the known mutual information $I_{\gamma}(X;Y)$ as the label. The value of $n$ varies between $10$ and $500$ across each of our train and test meta-datasets, thus focusing on low-sample regimes. The resulting training meta-dataset $\mathcal{M}_{\text{train}}$ contains 16 base distribution families and 625,000 meta-datapoints. $\mathcal{M}_{\text{test-extended}}$ contains 806,000 meta-datapoints from 13 distribution families, of which 6 are shared with $\mathcal{M}_{\text{train}}$. Meanwhile, $\mathcal{M}_{\text{test}}$ has a reduced size of 2,340 meta-datapoints, created to allow the testing of slower baselines that would take prohibitively long to run on the extended test set. The test sets are sampled to uniformly cover all sample sizes, dimensions, and base distribution families. The contents and generation of the meta-datasets are described in more detail in Appendix B.
Model Architecture
Building an estimator using our framework presents three main architectural challenges. First, the model must process datasets of varying size while remaining invariant to sample ordering. The Transformer architecture is naturally permutation-invariant and has been a common choice to ingest datasets as inputs (transformer_invariance). The SetTransformer (set_transformer) architecture introduced the ISAB block to reduce the quadratic cost of attention by performing attention on a fixed number of learned inducing points. Finally, the SetTransformer++ architecture (set_transformer_plus_plus) further improves upon the original SetTransformer by introducing set normalization and additional residual connections that enable more stable training with deeper models. We therefore use SetTransformer++ as the basis of our architecture. Second, our estimator must process samples of varying dimensions. To address this, we introduce an additional attention block that operates over the dimension axis; this time, permutation invariance is not required. This row-wide attention also operates with a learned fixed inducing point, which then serves as a pooling mechanism into a fixed row dimension. Third, unlike many supervised learning tasks, mutual information is an unbounded measure. This is a long-standing challenge for MI estimators, which typically underestimate MI and fail in high-MI settings. We explored several normalization and multi-task prediction strategies (e.g., predicting normalized conditional entropies) to address this problem when designing our model, and provide our results in Appendix C.1. We found that the simplest solution of directly predicting MI produced the best estimators, provided the meta-dataset includes a sufficient range of MI labels.
4 Experiments
Empirically, we find that our learned estimators outperform existing methods across sample sizes and dataset dimensions, particularly in low-sample settings (4.1). We provide an empirical study of the behavior of estimators under increasing dimensionality, finding that on average, our models require roughly half the number of samples as the best baseline to provide reliable MI estimates (4.2). The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives (4.3), while inference is orders of magnitude faster than existing neural methods (4.4). Finally, the learned estimators generalize to unseen distributions and sample sizes (4.5), and benefits from the mix of dimensions and size of the training dataset (4.6).
4.1 Comparison with Existing Estimators
We begin by benchmarking our estimators against existing methods on a diverse set of synthetic distributions. In Table 1, we present the average MSE loss of each estimator on $\mathcal{M}_{\text{test}}$, grouped by ranges of meta-datapoint size for clarity. We use the smaller test set of 2,340 meta-datapoints to evaluate all baselines.
In-Meta-Distribution (IMD) Out-of-Meta-Distribution (OoMD)
$n\in$ $[10,100]$ $[100,300]$ $[300,500]$ $[10,100]$ $[100,300]$ $[300,500]$
CCA $4.6\mathrm{e}^{3}$ $\pm 7.4\mathrm{e}^{3}$ $2.3\mathrm{e}^{1}$ $\pm 7.1\mathrm{e}^{1}$ $6.7\mathrm{e}^{0}$ $\pm 3.7\mathrm{e}^{1}$ $4.0\mathrm{e}^{3}$ $\pm 6.8\mathrm{e}^{3}$ $1.8\mathrm{e}^{1}$ $\pm 7.3\mathrm{e}^{1}$ $1.2\mathrm{e}^{1}$ $\pm 6.2\mathrm{e}^{1}$
KSG $3.0\mathrm{e}^{1}$ $\pm 8.7\mathrm{e}^{1}$ $2.8\mathrm{e}^{1}$ $\pm 1.0\mathrm{e}^{2}$ $2.7\mathrm{e}^{1}$ $\pm 9.5\mathrm{e}^{1}$ $3.7\mathrm{e}^{1}$ $\pm 7.7\mathrm{e}^{1}$ $3.5\mathrm{e}^{1}$ $\pm 8.6\mathrm{e}^{1}$ $4.3\mathrm{e}^{1}$ $\pm 1.1\mathrm{e}^{2}$
MINE $6.3\mathrm{e}^{3}$ $\pm 3.7\mathrm{e}^{4}$ $1.2\mathrm{e}^{4}$ $\pm 1.2\mathrm{e}^{5}$ $1.2\mathrm{e}^{5}$ $\pm 2.0\mathrm{e}^{6}$ $7.1\mathrm{e}^{3}$ $\pm 4.1\mathrm{e}^{4}$ $1.5\mathrm{e}^{4}$ $\pm 2.3\mathrm{e}^{5}$ $3.3\mathrm{e}^{4}$ $\pm 4.0\mathrm{e}^{5}$
InfoNCE $1.4\mathrm{e}^{2}$ $\pm 2.5\mathrm{e}^{2}$ $3.5\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$ $3.2\mathrm{e}^{1}$ $\pm 1.1\mathrm{e}^{2}$ $3.2\mathrm{e}^{2}$ $\pm 2.1\mathrm{e}^{3}$ $4.2\mathrm{e}^{1}$ $\pm 9.7\mathrm{e}^{1}$ $5.0\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$
DV $3.0\mathrm{e}^{19}$ $\pm 2.1\mathrm{e}^{20}$ $1.1\mathrm{e}^{9}$ $\pm 2.2\mathrm{e}^{10}$ $8.5\mathrm{e}^{6}$ $\pm 6.7\mathrm{e}^{7}$ $1.2\mathrm{e}^{19}$ $\pm 6.3\mathrm{e}^{19}$ $8.7\mathrm{e}^{6}$ $\pm 5.1\mathrm{e}^{7}$ $6.1\mathrm{e}^{6}$ $\pm 5.8\mathrm{e}^{7}$
LMI $3.1\mathrm{e}^{1}$ $\pm 9.1\mathrm{e}^{1}$ $3.1\mathrm{e}^{1}$ $\pm 1.1\mathrm{e}^{2}$ $3.0\mathrm{e}^{1}$ $\pm 1.1\mathrm{e}^{2}$ $3.8\mathrm{e}^{1}$ $\pm 8.0\mathrm{e}^{1}$ $3.6\mathrm{e}^{1}$ $\pm 9.1\mathrm{e}^{1}$ $4.6\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$
MINDE $3.5\mathrm{e}^{1}$ $\pm 9.6\mathrm{e}^{1}$ $3.2\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$ $2.5\mathrm{e}^{1}$ $\pm 9.1\mathrm{e}^{1}$ $4.2\mathrm{e}^{1}$ $\pm 8.5\mathrm{e}^{1}$ $3.8\mathrm{e}^{1}$ $\pm 9.4\mathrm{e}^{1}$ $3.6\mathrm{e}^{1}$ $\pm 1.0\mathrm{e}^{2}$
MIST $3.1\mathrm{e}^{0}$ ± $10.4$ $3.4\mathrm{e}^{-1}$ ± $0.9$ $2.8\mathrm{e}^{-1}$ ± $0.9$ $8.0\mathrm{e}^{0}$ ± $14.1$ $7.3\mathrm{e}^{0}$ ± $18.5$ $8.8\mathrm{e}^{0}$ ± $22.3$
MISTQR $3.1\mathrm{e}^{0}$ ± $11.7$ $3.1\mathrm{e}^{-1}$ ± $1.5$ $1.2\mathrm{e}^{-1}$ ± $0.3$ $9.5\mathrm{e}^{0}$ ± $17.1$ $9.9\mathrm{e}^{0}$ ± $24.6$ $1.0\mathrm{e}^{1}$ ± $25.4$
Table 1: Comparison of MI estimation baselines and our proposed methods across different distribution families (IMD and OoMD) and varying numbers of samples per meta-datapoint on the $\mathcal{M}_{\text{test}}$. Our estimators substantially outperform existing methods across sample size and distribution settings, including for distributions unseen during training. Shown are the averaged MSE loss values (mean ± stddev) of each estimator, grouped by the number of samples $n$ given to the estimator. The standard deviation is computed from 100 bootstrap resamples for each regime.
In this challenging regime of low sample sizes, varying dimensionality, and non-Gaussian distributions, recent estimators struggle to outperform the classical KSG baseline. This aligns with recent findings demonstrating that evaluation on more complex settings reveals weaknesses in modern estimators (bmi_benchmark; lee2024a).
Our learned estimators substantially outperform all baselines. Relative to KSG (the next strongest method), our models achieve approximately $10\times$ lower error on distributions seen during training (IMD), and $\sim 5\times$ lower error on unseen distributions (OoMD). A breakdown by distribution family is provided in Appendix C.3. For subsequent large-scale evaluation, we focus on KSG as the baseline, since it is both computationally feasible on the extended test set and remains the strongest baseline estimator in Table 1, particularly in low-sample regimes.
[…]
Figure 2: Heatmaps providing a detailed analysis of the performance of the two proposed MI estimators and the KSG baseline. The data dimensions are grouped into three categories, and within each group, meta-datapoints are sorted by the number of samples. Shown are the average values of MSE, bias, variance, and confidence interval (CI cov.) width across $\mathcal{M}_{\text{test-extended}}$. CI coverage denotes the fraction of samples per group for which the true MI lies within the 95% bootstrap confidence interval.
[…]
Figure 3: Predicted MI as a function of the true MI on $\mathcal{M}_{\text{test-extended}}$. For each estimator, predictions are aggregated into bins of true MI values.
One of the most well-documented failure modes of existing MI estimators is their negative bias, in which estimators substantially underestimate predictions for distributions with large MI. In Figure 3, we plot the predictions of our estimators in terms of increasing MI. While KSG’s estimates plateau quickly, our estimates closely follow the true MI values.
4.2 Scaling Behavior
Traditional evaluations of MI estimators have been constrained to relatively simple distributions and high-sample regimes. While recent efforts such as the BMI benchmark have enabled more comprehensive studies on complex distributions, systematic empirical analysis of scaling behavior has remained limited, largely due to the computational expense of evaluating existing methods. Leveraging our larger test set $\mathcal{M}_{\text{test-extended}}$, we conduct a detailed empirical study of estimator performance across a wide range of dimensions and sample sizes. This enables us to characterize performance frontiers in terms of data requirements. This analysis addresses an important question for practitioners: How many samples are needed for a reliable estimate?
Figure 4 shows the data requirement profiles of our estimators compared to the KSG estimator. The x-axis represents the dimensionality of the distribution. The y-axis indicates the sample size required to achieve target MSE thresholds of $0.03$, $0.07$, and $0.09$, selected based on our prior heatmap analysis. Our estimators consistently attain the desired accuracy with substantially fewer samples than KSG. In particular, for certain MSE targets, even a sample size of approximately 500 is insufficient for KSG to achieve reliable performance. A complete heatmap of these results is provided in Appendix C.5.
[…]
4.3 Calibration
[…]
4.4 Inference Efficiency
[…]
We estimated the inference time of each baseline on the entirety of $\mathcal{M}_{\text{test-extended}}$ and report the results in Appendix C.4. The high computational costs (>> 1 week) of evaluating all baselines except for CCA and KSG on $\mathcal{M}_{\text{test-extended}}$, along with the poor performance measured for CCA on the $\mathcal{M}_{\text{test}}$ for low sample sizes, further motivate our selection of KSG as the only baseline for other experiments in this work.
Figure 6: Inference time for each method as a function of the sample size. Results are averaged over 100 inferences (5 for MINDE), each performed over 10 meta-datapoints.
4.5 Generalization to Unseen Distributions, Dimensions, and Sample Sizes
The previous sections demonstrate that the learned algorithm generalizes effectively to unseen distribution families, which is evaluated at dimensions and sample sizes seen during training. We find that by learning to infer MI across diverse distributions and sample regimes, our estimators learn appropriate regularization strategies that transfer even to distribution families not encountered during training. The strength of these results motivates a bolder question: if meta-learned MI algorithms can generalize to unseen distributions, can they also generalize to these distributions at dimensions and sample sizes beyond their training experience?
We conduct a controlled study by training models on restricted subsets of the training data, specifically training on meta-datapoints in $M_{\text{train}}$ with a dimensionality $D<16$ and a sample size $S<300$. We evaluate performance on our full test meta-dataset $\mathcal{M}_{\text{test-extended}}$, which therefore includes distributions with higher dimensionalities and sample sizes and new distributions. Figure 7 presents two loss heatmaps, one for $\mathcal{M}_{\text{test-extended,IMD}}$ and one for $\mathcal{M}_{\text{test-extended,OoMD}}$.
We find that our estimators generalize well to new dimensions and sample sizes for distributions that are in the training meta-distribution (IMD heatmap). Similarly, we find that our models generalize well to new distributions when the data dimensionality is similar to that during training (OoMD heatmap, left half). However, they struggle to generalize to unseen distributions when the data also has higher dimensionality than any seen during training (OoMD heatmap, right half).
[…]
4.6 Impact of Variable Input Dimensions
A natural question that arises from our approach is the impact of including multiple dimensions in model training. Of course, training on a mix of dimensions allows the development of a general-purpose estimator that can achieve high performance across settings, as shown in the above experiments. However, there may be settings where practitioners know their data dimension apriori, and may consider producing a training dataset of only that dimension. To study these questions, we trained separate (specialized) copies of MIST on subsets of $\mathcal{M}_{\text{train}}$ filtered to contain only meta-datapoints with a fixed input dimension (e.g., 4, 8, 16 or 32). This filtering procedure reduces the number of training data points substantially (by a factor of 31). To provide a comparison with the same number of samples, we also trained an additional (reduced data) MIST model on a uniformly down-sampled version of $\mathcal{M}_{\text{train}}$ to match the training size of the individual dimension models. By comparing the base MIST (full training set) with the specialized models (only one dimension), we can measure the impact of scaling both the dimensions and dataset size. By comparing the reduced data model (downsampled train set) and the specialized models (only one dimension), we can isolate the impact of training on multiple dimensions, without increasing dataset scale. In Figure 8, we provide these comparisons for each input dimension in $\mathcal{M}_{\text{test-extended}}$.
[…]
5 Discussion
This work rethinks the design of mutual information estimators. Rather than deriving estimators analytically from variational or density-based formulations, we treat estimation as a supervised learning problem: the estimator is a learned function, trained end-to-end on a large synthetic meta-dataset with known ground-truth MI values. This approach represents a deliberate shift in paradigm, trading theoretical motivations for a flexible and data-driven design.
Trade-offs and limitations. By embracing an empirical learning framework, we inherit the usual risks of statistical learning. The learned estimator’s accuracy depends on the diversity and coverage of the meta-training distributions, and its generalization to distant out-of-distribution families cannot be guaranteed a priori. Our estimator’s behavior is learned, rather than designed, and therefore does not carry the theoretical consistency guarantees of classical MI estimators—namely, that the estimate converges to the true MI as $n\to\infty$. We studied our estimator’s behavior on self-consistency tests from MINE_exp_variance_consistency_tests in Appendix D, and find that while our estimators are approximately self-consistent under the independence test, they are not self-consistent and produce high variance prediction ratios for the data processing and additivity tests. We believe that a learning-based MI framework offers opportunity for designing models, datasets, and losses to induce such desired behavior characteristics, but we believe it is important to recognize that our estimators do not carry these properties innately. In high-sample settings, our learned estimator cannot directly leverage all available samples during one inference pass and other estimators may be preferred if time and compute resources are available. Moreover, learned estimators are not inherently interpretable and may result in unforeseen failure modes. These limitations parallel the general challenges faced by machine learning models in other domains, highlighting the importance of careful validation and uncertainty quantification.
Controllability and adaptability. A key strength of this framework lies in its controllability. Our estimators’ behavior are largely determined by the synthetic datasets created; practitioners can customize, interrogate, and design datasets to improve estimator quality and reliability. Since mutual information is invariant under invertible transformations, our meta-dataset can be mapped to a target modality using normalizing flows without altering its ground-truth MI. This property enables the generation of diverse synthetic meta-datasets that reflect the characteristics of arbitrary data domains, such as images, text, tabular data, or multimodal signals. By tailoring the training data through such transformations, one can design estimators specialized for particular classes of distributions or downstream applications. This mechanism opens the door to what may be viewed as a foundational model for information-theoretic estimation, one that can be adapted via fine-tuning to specific tasks and modalities.
Computational efficiency. Once trained, our learned estimators perform inference in a single forward pass (or two to obtain a confidence interval using the quantile regression model), making them several orders of magnitude faster than existing neural estimators. By amortizing the cost of inference during meta-training, our estimators achieve speedups of three to four orders of magnitude compared to neural MI estimators, and one order of magnitude compared to classical KSG implementations. The efficiency of our estimator offers new opportunities for MI estimation in settings where existing estimators are infeasible or intractable.
Outlook. Beyond empirical gains, the proposed framework opens numerous directions for research in both information theory and statistical learning. First, it allows the inclusion of theoretical results as inductive biases, for instance, explicitly penalizing violations of the Data Processing Inequality. Second, because MI estimators are often used within optimization loops (e.g., to maximize or minimize information), it is possible to train models that focus on relative ordering rather than absolute accuracy, for example, by using ranking losses that emphasize correct ordering across a batch of datasets. Taken together, these elements outline a research program: the development of trainable, meta-learned estimators as flexible, efficient, and customizable alternatives to classical estimators.
Acknowledgments
This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2039655 (MR). Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors(s) and do not necessarily reflect the views of the National Science Foundation.
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) with a grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research. This work was also supported by the Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI) and the National Science Foundation (under NSF Award 1922658).
This work was partially conducted within French research unit UMR 5217 and was supported by CNRS (grant ANR-22-CPJ2-0036-01) and by MIAI@Grenoble-Alpes (grant ANR-19-P3IA-0003). It was granted access to the HPC resources of IDRIS under the allocation 2025-AD011014834 made by GENCI.
Appendix A Appendix
A.1 Definitions of Mutual Information
Several equivalent expressions of mutual information are leveraged by existing estimators. For clarity, we provide a simple overview of these expressions below. For two continuous random variables $X$ and $Y$, the mutual information (MI) is defined in terms of the joint distribution $P_{(X,Y)}$ and marginal distributions $P_{X}$ and $P_{Y}$:
$I(X;Y)=\int_{\mathcal{Y}}\int_{\mathcal{X}}P_{(X,Y)}(x,y)\log\left[\frac{P_{(X,Y)}(x,y)}{P_{X}(x)P_{Y}(y)}\right]dx\,dy$ (4)
Donsker-Varadhan
A commonly used expression of MI is in terms of the Kullback-Leibler (KL) divergence between the joint distribution and the product of marginals (donsker_varadhan_kl_div_mi_definition):
$I(X;Y)=D_{\text{KL}}(P_{(X,Y)}\|P_{X}\otimes P_{Y})$ (5)
This expression can be understood as measuring by how much the joint distribution deviates from the product of marginal distributions. If $X$ and $Y$ are independent, $P_{(X,Y)}=P_{X}\otimes P_{Y}$, producing a KL-divergence (and MI estimate) of zero.
Conditional Entropy
Another common expression of MI is in terms of the conditional entropy: $H(\cdot|\cdot)$ denotes conditional entropy, and $H(X,Y)$ is the joint entropy.
I(X;Y)=H(X)H(X|Y)=H(Y)H(Y|X)=H(X)+H(Y)H(X,Y)I(X;Y)=H(X)-H(X|Y)=H(Y)-H(Y|X)=H(X)+H(Y)-H(X,Y) (6)
This expression of MI can be understood as a measure of the reduction in uncertainty about one variable given the other.
A.2 MI Estimators
We provide an overview of prominent MI estimators for reference in Table 2, categorized by their prediction category, the MI definition considered, and the estimation task. Previous works such as MINE_exp_variance_consistency_tests have described variational estimators as either discriminative or generative, with discriminative methods learning the ratio between joint and marginal densities directly and generative methods estimating the densities separately before computing the density ratio. We broaden this categorization into density and density ratio estimators.
Method Category $I(X;Y)$ Estimation Task
Binning Density $\sum_{x,y}p(x,y)\log\frac{p(x,y)}{p(x)p(y)}$ Estimate $p(x,y)$, $p(x)$, $p(y)$ via histograms
KSG Density =H(X)+H(Y)H(X,Y)=H(X)+H(Y)-H(X,Y) Implicitly estimate H()H(*) by writing H()H(*) in terms of KNN distances
KDE Density $\sum_{x,y}p(x,y)\log\frac{p(x,y)}{p(x)p(y)}$ Estimate $p(x,y)$, $p(x)$, $p(y)$ via kernels
MIENF Density $I(f_{X}(X);f_{Y}(Y))$ Learn flows $f_{X}$, $f_{Y}$ to Gaussian marginals
LMI Density $\approx I(Z_{x};Z_{y})$ Learn neural networks for embeddings $Z_{x}=g_{X}(X)$, $Z_{y}=g_{Y}(Y)$, then apply KSG
GM Density $\int\int p(x,y)\log\frac{p(x,y)}{p(x)p(y)}dx\,dy$ Estimate $p(x,y)$, $p(x)$, $p(y)$ via VAEs (or flows)
MINDE Density $D_{KL}(p(x,y)\|p(x)p(y))$ Learn scores $\nabla\log p(x,y)$, $\nabla\log p(x)$ $\nabla\log p(y)$
VCE Density $H[c(\mathbf{u}_{X},\mathbf{u}_{Y})]$ Learn marginal ranks $\mathbf{u}_{X}$, $\mathbf{u}_{Y}$ with flows, estimate vector copula $c$ with MLE over mixture of Gaussian copulas
MINE Density ratio $\geq\sup_{\theta}\mathbb{E}_{P_{XY}}[T_{\theta}]-\log(\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}}])$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approx. $\log(\frac{p(x,y)}{p(x)p(y)})$
NWJ Density ratio $\geq\sup_{\theta}\mathbb{E}_{P_{XY}}[T_{\theta}]-\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}-1}]$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approx. $\log\frac{p(x,y)}{p(x)p(y)}$
InfoNCE Density ratio $\geq\mathbb{E}_{(x,y)\sim P_{n}}\left[\log\frac{f_{\theta}(x,y)}{\frac{1}{n}\sum_{j=1}^{n}f_{\theta}(x,y_{j})}\right]$ Learn classifier critic $f_{\theta}(x,y)\propto\frac{p(x,y)}{p(x)p(y)}$
SMILE Density ratio $\geq\mathbb{E}_{P}[T_{\theta}]-\log\mathbb{E}_{Q}[\text{clip}{\tau}(e^{T_{\theta}})]$ * Learn classifier critic $f_{\theta}(x,y)$ to maximize bound.
MIST From Samples $\approx F(x,y)$ Learn estimator $F$ by direct optimization
Table 2: Summary of prominent methods for estimating mutual information, along with the MI definition used and their estimation tasks.
$\text{clip}_{\tau}(z)\in[e^{-\tau},e^{\tau}]$.
A.3 Empirical Evaluation Scope of MI Estimators
In Table 3, we provide a summary of how prominent estimators were empirically evaluated. We focus our summary on evaluations of the estimator directly, rather than the performance of an estimator in a downstream tasks. Similarly, we only include the evaluations of estimators in the papers introducing the method, as we are most interested in how estimator quality is determined.
Method Data Type # $\mathcal{D}$ # $\not\sim\mathcal{N}$ $d_{\max}$ $I_{\max}$ $n_{\min}$
Binning Time-Series 1 1 1 - 65K
KDE Time-Series 4 4 1 - 400, 2K, 4K
KSG Synthetic 4 3 3, 8 0.83,1,1.2 125, 10K, 50K
MIENF Synthetic 6 5 32, 1024 10 10K
LMI Synthetic,MNIST,PLM 3 2 5K,784,1024 1 2K,5K
GM Synthetic 2 1 20 10 256K
MINDE Synthetic 20, 3 17, 2 25, 9 1, 5 100K
VCE Synthetic,Images,Text 7,2,2 6,1,2 64,16,16 16,7,2.1 10K,10K,4K
MINE Synthetic 2 0 2, 20 2, 40 -
NWJ - - - - - -
InfoNCE - - - - - -
SMILE Synthetic 2 1 20 10 256K
MIST Synthetic 23 19 32 40 10
Table 3: Summary of the empirical evaluation scope for prominent mutual information estimators. Comma-separated entries indicate values for different experimental settings (e.g., KDE has three experiments with minimum sample size considered 400, 2K, and 4K, respectively). Dashes indicate data that is not available or not specified.
Early methods for mutual information estimation, such as binning, were evaluated primarily using simulated time-series data. These commonly included sine and cosine functions, as well as chaotic dynamical systems such as the Lorenz and Rossler attractors (where the task is phase reconstruction, and no ground truth is available). In the initial histogram binning work (histogram_binning), the Rossler attractor was used. For this work, we include it in the table as having one 1D distribution, which is non-Gaussian, with no ground truth MI reported (and therefore no maximum). The experiment calculated the mutual information over 65,536 points.
KDE (kde) was similarly evaluated with time-series data. KDE was evaluated on 1D synthetic sine waves, an autoregressive function of a sine and a Gaussian distribution, the Lorenz attractor, and the Rossler attractor. 400 samples were used for the sine function, and 500 samples were used for the autoregressive function. 4096 and 2048 samples were used for the Lorenz and Rossler attractors, respectively. Estimators are evaluated on the quality of the reconstruction, with no ground truth MI available. We include KDE in the table as having 4 distributions (sine, AR, Lorenz, Rossler), 4 of which are non-Gaussian, and all 1D. No ground truth MI values are present, and the samples included 400, 2K, 4K.
KSG (ksg) performed evaluation on 2D correlated Gaussians with unit variance and zero mean, testing sample sizes of 125, 250, 500, 1K, 2K, 4K, 10K, and 20K. The MI was 0.830366. Authors performed error analysis with 10K samples, and performed experiments on the gamma-exponential distribution, the ordered Weinman exponential, and the circle distribution. All of these had MI values below 1, and considered samples with 1K and 10K data points. For their higher dimension experiments, the estimator was evaluated on 3-D Gaussian distributions with a sample size of 10K in one experiment, and on m-correlated uniform Gaussians up to dimension 8 with a sample size of 50K. We include KSG in the table as being evaluated on synthetic 4 distributions (Gaussian, Weinman-exponential, gamma-exponential, and circle), 3 of which are non-Gaussian. The maximum dimension considered was 8, while the maximum MI values were 0.83, 1, and 1.2 for the 1-D Gaussians, 3-D Gaussians, and 8-D Gaussian experiments, respectively. The minimum sample size in these experiments was 125 for the 1-D Gaussian, 10K for other distributions, and 50K for larger dimensions.
MIENF (mi_through_normalizing_flows) first evaluated their estimator on 2D correlated Gaussian distributions mapped into higher-dimensional images (16x16 and 32x32), evaluated on 10K samples, and with MI values between 0 and 10. In their second experiment, authors evaluated their method on five non-compressible distributions (uniform, smoothed uniform, student (dof = 4), arcsignh (student) student with dof = 1,2). For this experiment, authors varied the dimensions of these distribution from 2 to 32, using 10K samples for each prediction. The MI was set to 10 for all distributions and dimensions.
Latent-MI (LMI; lmi_estimator) was evaluated on multivariate Gaussian distributions under varying original data dimensionality and latent dimensionality learned before applying KSG. Their multivariate Gaussian experiment considered up to 10 latent dimensions and 5K original dimensions, with 2K samples. In a second experiment, the authors studied the sample requirements of their estimators. For this experiment, authors used multivariate Gaussians from dimension 1 to 50, each with one correlated dimension, and an MI of 1. Authors varied the sample size up to 4K. In a third experiment, MNIST and protein embeddings were resampled to align with a Bernoulli vector with set correlation. For MNIST, the dimension was 784, and the sample size was 5K. For the protein embeddings, the dimension was 1024 and the sample size was 4402. In both cases, datasets were generated with MI values between 0 and 1.
For GM and SMILE (MINE_exp_variance_consistency_tests), evaluation was performed on two tasks. In the first task, authors evaluated their methods with 20-D Gaussians with correlation $p$, and in the second task, the authors applied a cubic operation on these Gaussians. This set-up was introduced in (mine). MI values were increased from 0 to 10 at step intervals during the critic training process. Training occurred over 20K steps, with MI increasing by 2 every 4K steps. The batch size was 64, and each batch was a new random sample. For the sake of our table, we consider the number of samples used for estimation to be the number of samples used to fit the critic for a given MI value. Since the estimator saw 4K steps of batch size 64 for each MI (2, 4, 6, 8, 10), we list the minimum samples seen as 4,000 x 64 = 256K. This work also includes consistency tests constructed by adapting MNIST images, but this setting does not carry ground-truth MI values (only the ideal ratio of estimates).
MINDE (minde) used the BMI benchmark (bmi_benchmark) at larger sample sizes (100K rather than 10K). For our table, we consider the number of unique distribution types considered; for BMI, there are 20 unique distribution tasks after excluding tasks that are higher-dimensional versions of another. The highest MI value used was 5, and the maximum dimension considered was 25. MINDE also performs image consistency experiments as in SMILE, but this setting does not have a ground truth value.
VCE chen2025neural_vector_copulas evaluate estimators on five synthetic distributions with varying dependence. In this setting, MI ranges from 0 to 16, 10K samples are used, and dimensions range from 2 to 64. Then, the authors study increasing dimensionality on five distributions, with dimensions from 2 to 64 and MI values up to 14. 10K samples were used, and six distributions across the two synthetic experiments were not gaussian. Next, authors consider correlated rectangles and gaussian plates. These images have dimension 16x16 but are processed by an autoencoder prior to prediction to produce 16 dimensional embeddings. The max MI for this setting is 7 and 10K samples were used. Finally, they consider language embeddings. Data is similarly processed by an autoencoder to be 16 dimensions. The maximum MI is 2.1, and 400 samples were used.
MINE (mine) was first evaluated on 2-D and 20-D Gaussians with MI from 0 to 2 and 0 to 40, respectively. The number of samples used was not specified. In a second experiment, the authors applied a deterministic nonlinear transformation and demonstrated that their estimates are invariant to it.
NWJ (nwj) introduced a variational bound for estimating the KL-divergence between two arbitrary distributions, and their experiments focused on evaluating the quality of this bound and estimation on various distributions. The authors did not evaluate this approach on mutual information prediction directly, for which the distributions compared would be a joint distribution and the product of its marginals.
InfoNCE (oord2018_cpc) was developed as a method for representation learning, with evaluation including representation learning tasks but without any evaluation of MI prediction itself.
Appendix B Details of Meta-Dataset Creation
We construct a meta-dataset of synthetic distributions with known MI using the BMI library (bmi_benchmark), which generates complex yet tractable distributions via invertible transformations applied to base families with analytical MI. Each meta-sample consists of joint samples from a distribution paired with its ground-truth MI value, enabling supervised learning directly over distributions.
To promote diversity and generalization, we partition distribution families into disjoint training $\mathcal{D}_{\text{train}}$ and testing $\mathcal{D}_{\text{test}}$ groups and vary the sample dimensionality from $2$ to $32$. Since many existing estimators are prohibitively slow at this scale, we construct two evaluation corpora: a smaller $\mathcal{M}_{\text{test}}$ benchmark set for comparison with existing methods, and a larger $\mathcal{M}_{\text{test-extended}}$ set for assessing generalization to novel and higher-dimensional distributions. Complete details of meta-dataset sizes and included distribution families are provided in Table 4.
The dataset includes both In-Meta-Distribution (IMD) and Out-of-Meta-Distribution (OoMD) families, differing in their base distributions and MI-preserving transformations.
In-Meta-Distribution (IMD)
The IMD test set comprises samples drawn from the following base distributions, each with its own structured variants and associated sampled hyperparameters:
  • multi_normal: Multivariate normal distributions over $(X,Y)$ with jointly Gaussian structure and identity marginal variances. Structured variants are parameterized as follows:
    • dense: off_diag $\sim\mathcal{U}(0.0,0.5)$ (uniform pairwise correlation between all variables)
    • lvm: n_interacting $\in\{1,\dots,5\}$ (latent dimensionality)
      alpha $\sim\mathcal{U}(0.0,1.0)$ (anisotropy; relative scaling of $X$ vs. $Y$ loadings)
      lambd $\sim\mathcal{U}(0.1,3.0)$ (latent signal strength)
      beta $\sim\mathcal{U}(0.0,1.5)$ (additive noise scale in $X$)
      eta $\sim\mathcal{U}(0.1,5.0)$ (intensity of nonlinear warping in $X$)
  • multi_student: Multivariate Student’s $t$-distributions over $(X,Y)$ sharing the same covariance (or scatter) structure as their Gaussian counterparts, but with heavier tails controlled by the degrees of freedom. Structured variants are parameterized as follows:
    • dense: off_diag $\sim\mathcal{U}(0.0,0.5)$ (uniform pairwise correlation)
      df $\sim\mathcal{U}(1,10)$ (degrees of freedom; lower values correspond to heavier tails)
    • sparse: n_interacting $\in\{1,\dots,5\}$ (latent dimensionality; governs low-rank structure)
      strength $\sim\mathcal{U}(0.1,5.0)$ (latent signal strength; used as lambd and eta)
      df $\sim\mathcal{U}(1,10)$ (degrees of freedom)
Out-of-Meta-Distribution (OoMD)
The OoMD test set comprises distribution families that are absent from the training set. It focuses on the following base distribution:
  • multi_additive_noise: A multivariate additive noise model in which, for each dimension $j=1,\dots,d$,
    $X_{j}\sim\mathrm{Uniform}(0,1),\quad N_{j}\sim\mathrm{Uniform}(-\epsilon,\epsilon),\quad Y_{j}=X_{j}+N_{j},$
    with a shared noise scale $\epsilon\sim\mathcal{U}(0.1,2.0)$. This distribution is non-Gaussian with bounded support, and its mutual information equals the sum of per-dimension contributions.
    Sampled hyperparameter:
    • epsilon $\sim\mathcal{U}(0.1,2.0)$ (half-width of the uniform noise interval; controls the signal-to-noise ratio)
The true MI values for these distributions and their transformations are computed analytically or numerically, depending on tractability.
Dataset Split Size (Extended / General) Distribution Families
$\mathcal{M}_{\text{train}}$ 624652 multi_normal-dense-[base, asinh, halfcube],
multi_normal-lvm-[base, wigglify],
multi_normal-sparse-[base, wigglify],
multi_normal-2pair-base,
multi_student-dense-[base, wigglify],
multi_student-sparse-[base, asinh, halfcube],
multi_student-2pair-[asinh, halfcube],
multi_additive_noise-wigglify
$\mathcal{M}_{\text{test, IMD}}$ 372000 / 1080 multi_normal-dense-[base, halfcube],
multi_normal-lvm-wigglify,
multi_student-dense-base,
multi_student-sparse-[asinh, base]
$\mathcal{M}_{\text{test, OoMD}}$ 434000 / 1260 multi_normal-dense-wigglify,
multi_normal-sparse-[asinh, halfcube],
multi_student-dense-halfcube,
multi_student-sparse-wigglify,
multi_additive_noise-[base, halfcube]
Table 4: Meta-Dataset Description. Each distribution family is labeled as multi_<base_dist>-<structure>-<transform>, where <base_dist> denotes the base distribution type, <structure> specifies the parameterization structure, and <transform> indicates the nonlinear transformation applied to the data. For resource-constrained evaluation, a reduced test subset of 1k samples is used for both IMD and OoMD splits. The reduced subsets preserve the same distribution family composition as their full counterparts.
Appendix C Results
C.1 Experiments for Normalized and Multi-Task Predictions
In high-dimensional settings, estimated MI values increased substantially, often reaching several dozen, as illustrated in Figure 10. Such a wide numerical range can make direct regression with MSE or QCQR losses unstable, since large target values can lead to unstable gradients and slow convergence. To mitigate this issue, we applied normalization to the target variable, constraining its range and improving numerical stability.
To stabilize regression over wide MI ranges, we evaluated two normalization schemes:
$\hat{I}_{\mathrm{log}}$ $=\log\!\big(I(X;Y)+1\big),$ $\hat{I}_{\mathrm{inv}}$ $=\frac{I(X;Y)}{I(X;Y)+1},$ (7)
which reduce the impact of outliers and transform regression into a constrained prediction problem.
In addition to target normalization, we enriched the model by predicting auxiliary entropy coefficients corresponding to $H(X)$, $H(Y)$, and $H(X,Y)$. Using the same monotonic mapping $f(z)=z/(1+z)$, the normalized forms were defined as:
$\hat{H}_{X,\mathrm{inv}}$ $=f\!\big(H(X)\big),$ $\hat{H}_{Y,\mathrm{inv}}$ $=f\!\big(H(Y)\big),$ $\hat{H}_{XY,\mathrm{inv}}$ $=f\!\big(H(X,Y)\big).$ (8)
We investigated two formulations: (i) predicting the three entropy coefficients and reconstructing MI via Eq. 6, and (ii) predicting all four coefficients $\big(\hat{I}_{\mathrm{inv}},\hat{H}_{X,\mathrm{inv}},\hat{H}_{Y,\mathrm{inv}},\hat{H}_{XY,\mathrm{inv}}\big)$ in the multi-task setting.
For the four-coefficient configuration, a physics-informed term was introduced to enforce consistency between two independent MI estimates. The first estimate is obtained from the directly predicted MI coefficient $\hat{I}_{1}=f^{-1}\!\big(\hat{I}_{\mathrm{inv}}\big)$ and the entropy-derived target $\hat{I}_{2}=f^{-1}\!\big(\hat{H}_{X,\mathrm{inv}}\big)+f^{-1}\!\big(\hat{H}_{Y,\mathrm{inv}}\big)-f^{-1}\!\big(\hat{H}_{XY,\mathrm{inv}}\big)$. To achieve alignment between these quantities, the overall training objective is expressed as
$\mathcal{L}_{\mathrm{total}}=\underbrace{\frac{1}{4}\sum_{j\in\{\mathrm{inv},X,Y,XY\}}\big\|\hat{Q}_{j}-y_{j}\big\|^{2}}_{\mathcal{L}_{\mathrm{base}}}\;+\;\lambda_{\mathrm{phys}}\underbrace{\mathbb{E}_{\mathcal{B}}\!\left[\left|\frac{\hat{I}_{1}}{\hat{I}_{2}+\epsilon}-1\right|\right]}_{\mathcal{L}_{\mathrm{phys}}},$ (9)
where $\hat{Q}_{j}\in\{\hat{I}_{\mathrm{inv}},\hat{H}_{X,\mathrm{inv}},\hat{H}_{Y,\mathrm{inv}},\hat{H}_{XY,\mathrm{inv}}\}$ denotes the predicted normalized coefficients, $\lambda_{\mathrm{phys}}=0.3$ balances regularization strength, $\epsilon>0$ prevents division by zero, and $\mathbb{E}_{\mathcal{B}}[\cdot]$ denotes expectation over a training batch. We trained this model on a fixed-dimension subset from $\mathcal{M}_{\text{train}}$ with all samples having $\mathrm{dim}=9$. The goal of this setup is to study whether predicting physically-related coefficients can help refine the model’s latent representations, incorporate additional structural knowledge, and improve prediction robustness.
Figure 9: Distribution of true MI values within $\mathcal{M}_{\text{train}}$.
Prediction Method $\mathcal{M}_{\text{test,IMD}}$ $\mathcal{M}_{\text{test,OoMD}}$
Direct MI $\mathbf{5.4\mathrm{e}^{-2}}$ $\pm 2.0$ $\mathbf{1.2\mathrm{e}^{0}}$ $\pm 2.3$
Log norm. ($\hat{I}_{\mathrm{log}}$) $8.5\mathrm{e}^{0}$ $\pm 2.1$ $1.4\mathrm{e}^{1}$ $\pm 2.4$
Inverse norm. ($\hat{I}_{\mathrm{inv}}$) $8.7\mathrm{e}^{-2}$ $\pm 2.5$ $1.9\mathrm{e}^{0}$ $\pm 3.3$
Entropy (3-coef) $2.5\mathrm{e}^{-1}$ $\pm 3.0$ $3.5\mathrm{e}^{0}$ $\pm 4.0$
Entropy & $\hat{I}_{\mathrm{inv}}$ (with $\mathcal{L}_{\mathrm{phys}}$) $\mathbf{5.2\mathrm{e}^{-2}}$ $\pm 1.8$ $\mathbf{1.5\mathrm{e}^{0}}$ $\pm 2.0$
Entropy & $\hat{I}_{\mathrm{inv}}$ (w/o $\mathcal{L}_{\mathrm{phys}}$) $7.0\mathrm{e}^{-2}$ $\pm 2.4$ $1.8\mathrm{e}^{0}$ $\pm 3.0$
Figure 10: Mean MI estimates across prediction strategies on $\mathcal{M}_{\text{test}}$ ($\mathrm{dim}=9$), comparing normalization methods and coefficient configurations. “3-coef” uses the estimators $(\hat{H}_{X,\mathrm{inv}},\hat{H}_{Y,\mathrm{inv}},\hat{H}_{XY,\mathrm{inv}})$; the last two entries show predictions with $\hat{I}_{\mathrm{inv}}$, with and without the physics-informed loss $\mathcal{L}_{\mathrm{phys}}$. Values are (mean ± stddev) over three runs.
Experiments on $\mathcal{M}_{\text{test}}$ with $\mathrm{dim}=9$ are summarized in Table 10. The best results were obtained using two methods: direct MI prediction, and the four-coefficient formulation with the physics-informed loss. The comparatively lower performance of the three-coefficient model can be attributed to error propagation, since each entropy term is predicted independently and their individual errors accumulate during MI reconstruction.
In this work, we proceed with the direct MI prediction strategy for subsequent experiments. Nevertheless, the four-coefficient formulation remains a promising direction for future research, as it provides richer structural information about the underlying distributions and could facilitate the derivation of both MI and entropy values within a unified framework.
C.2 Dimensionality Ablation
Figures 11 through 13 contain additional bar plots for the experiment described in the second half of Section 4.6.
Figure 11: Average MSE obtained by variable- and fixed-dimensionality versions of MIST on subsets of $\mathcal{M}_{\text{test-extended,OoMD}}$ with fixed dimensions.
[…]
We give in Figures 14 through 16 a high-level summary of the general performance of all various fixed- and variable-dimensionality models for ease of comparison.
Figure 14: Average MSE of the proposed models with variable input dimensionality versus dimension-specific variants. In this plot and the following ones, the test set was similar to the training one: variable-dimensionality models were evaluated on the full $\mathcal{M}_{\text{test-extended}}$, while other models were evaluated on subsets of it.
Figure 15: Average bias magnitude of the proposed models with variable input dimensionality versus dimension-specific variants.
Figure 16: Average variance of the proposed models with variable input dimensionality versus dimension-specific variants.
C.3 Comparing With Existing Estimators
[…]
CCA $7.7\mathrm{e}^{2}$ $1.1\mathrm{e}^{2}$ $2.0\mathrm{e}^{3}$ $1.8\mathrm{e}^{3}$ $1.3\mathrm{e}^{2}$ $9.1\mathrm{e}^{2}$ $1.2\mathrm{e}^{3}$ $1.5\mathrm{e}^{3}$ $8.8\mathrm{e}^{2}$ $1\mathrm{e}^{2} 1.1$$1\mathrm{e}^{3}$ $2.7\mathrm{e}^{2}$ $6.7\mathrm{e}^{2}$
KSG $3.2\mathrm{e}^{-2}$ $1.5\mathrm{e}^{1}$ $4.7\mathrm{e}^{-2}$ $1.6\mathrm{e}^{-1}$ $7.7\mathrm{e}^{1}$ $7.6\mathrm{e}^{1}$ $4.6\mathrm{e}^{-2}$ $5.7\mathrm{e}^{1}$ $7.5\mathrm{e}^{1}$ $5.2\mathrm{e}^{1}$ $2.5\mathrm{e}^{-1}$ $4.6\mathrm{e}^{1}$ $4.1\mathrm{e}^{1}$
MINE $1.2\mathrm{e}^{3}$ $1.8\mathrm{e}^{3}$ $1.8\mathrm{e}^{3}$ $2.3\mathrm{e}^{5}$ $7.8\mathrm{e}^{4}$ $4.3\mathrm{e}^{3}$ $2.8\mathrm{e}^{3}$ $1.8\mathrm{e}^{3}$ $2.1\mathrm{e}^{3}$ $6.7\mathrm{e}^{4}$ $6.9\mathrm{e}^{4}$ $5.4\mathrm{e}^{2}$ $3.5\mathrm{e}^{3}$
InfoNCE $1.5\mathrm{e}^{1}$ $3.9\mathrm{e}^{1}$ $1.4\mathrm{e}{{}^{1}}$ $2.0\mathrm{e}^{1}$ $1.1\mathrm{e}^{2}$ $1.3\mathrm{e}^{2}$ $1.3\mathrm{e}^{1}$ $8.0\mathrm{e}^{1}$ $1.1\mathrm{e}^{2}$ $3.2\mathrm{e}^{2}$ $3.2\mathrm{e}^{1}$ $7.1\mathrm{e}^{1}$ $6.9\mathrm{e}^{1}$
DV $1.4\mathrm{e}^{17}$ $2.1\mathrm{e}^{18}$ $6.2\mathrm{e}^{16}$ $3.9\mathrm{e}^{18}$ $6.5\mathrm{e}^{18}$ $2.3\mathrm{e}^{19}$ $1.5\mathrm{e}^{17}$ $1.8\mathrm{e}^{18}$ $2.3\mathrm{e}^{16}$ $1.6\mathrm{e}^{18}$ $6.0\mathrm{e}^{18}$ $6.0\mathrm{e}^{18}$ $1.6\mathrm{e}^{18}$
LMI $2.1\mathrm{e}^{-1}$ $1.6\mathrm{e}^{1}$ $1.4\mathrm{e}^{-1}$ $2.7\mathrm{e}^{-1}$ $8.3\mathrm{e}^{1}$ $8.4\mathrm{e}^{1}$ $1.9\mathrm{e}^{-1}$ $5.9\mathrm{e}^{1}$ $8.0\mathrm{e}^{1}$ $5.4\mathrm{e}^{1}$ $2.9\mathrm{e}^{-1}$ $4.7\mathrm{e}^{1}$ $4.2\mathrm{e}^{1}$
MINDE $8.8\mathrm{e}^{-1}$ $1.4\mathrm{e}^{1}$ $1.3\mathrm{e}^{0}$ $9.2\mathrm{e}^{-1}$ $8.2\mathrm{e}^{1}$ $8.0\mathrm{e}^{1}$ $6.8\mathrm{e}^{-1}$ $5.7\mathrm{e}^{1}$ $7.3\mathrm{e}^{1}$ $5.5\mathrm{e}^{1}$ $9.6\mathrm{e}^{-1}$ $4.4\mathrm{e}^{1}$ $3.6\mathrm{e}^{1}$
MIST $6.5\mathrm{e}^{-2}$ $1.5\mathrm{e}^{0}$ $5.4\mathrm{e}^{-2}$ $2.5\mathrm{e}^{-1}$ $1.4\mathrm{e}^{0}$ $2.0\mathrm{e}^{0}$ $1.6\mathrm{e}^{-1}$ $1.6\mathrm{e}^{0}$ $9.5\mathrm{e}^{-1}$ $2.0\mathrm{e}^{0}$ $3.4\mathrm{e}^{-1}$ $2.1\mathrm{e}^{1}$ $3.0\mathrm{e}^{1}$
MISTQR $3.4\mathrm{e}^{-2}$ $1.5\mathrm{e}^{0}$ $2.6\mathrm{e}^{-2}$ $1.4\mathrm{e}^{-1}$ $1.5\mathrm{e}^{0}$ $1.6\mathrm{e}^{0}$ $9.7\mathrm{e}^{-2}$ $1.4\mathrm{e}^{0}$ $6.9\mathrm{e}^{-1}$ $1.7\mathrm{e}^{0}$ $2.7\mathrm{e}^{-1}$ $3.0\mathrm{e}^{1}$ $3.6\mathrm{e}^{1}$
Multinormal (dense)
Wiggly @ Multinormal (lvm)
Half-cube @ Multinormal (dense)
Student-t (dense)
Student-t (sparse)
Asinh @ Student-t (sparse)
Wiggly @ Multinormal (dense)
Half-cube @ Multinormal (sparse)
Asinh @ Multinormal (sparse)
Wiggly @ Student-t (sparse)
Half-cube @ Student-t (dense)
Additive Noise
Half-cube @ Additive Noise
Table 5: Mean mutual information (MI) estimates across data distributions. Color indicates whether distributions are from the in-meta-distribution (blue) or out-of-meta-distribution (red) subsets from $\mathcal{M}_{\text{test}}$. We also highlight in green values that are within one order of magnitude of the best and mark the best values in bold.
C.4 Inference Time
We report in Table 6 the inference time of each of the baselines, averaged across the data from Figure 6.
Method Time / sample (s)
[…]
CCA 0.000 900.000\,90
KSG $0.021$
LMI $0.76$
InfoNCE $2.01$
NWJE $2.15$
DV $2.20$
MINE $2.75$
MINDE $122.7$
Table 6: Average inference time per sample measured on $\mathcal{M}_{\text{test-extended}}$.
C.5 Scaling Behavior
We report in Table 7 the performance of the main model (5.5M) used in our experiments compared to scaled-down models.
Model Params IMD OoMD
(M) MSE Abs. Bias Variance MSE Abs. Bias Variance
MIST 5.5 0.7508 3.0255 0.3401 7.6679 20.2033 0.3743
MISTQR 5.5 0.7637 0.6853 0.2941 9.2189 2.9803 0.3370
[…]
Table 7: Average performance metrics for models of different sizes.
We give in Figure 17 a fine-grained heatmap of the MSE performance obtained by MIST and MISTQR for various sample size groups and dimensions.
Figure 17: Detailed visualization of the average MSE values for our two proposed models across different sample-size groups and all evaluated dimensionalities. The measurements were obtained on $\mathcal{M}_{\text{test-extended,IMD}}$. The results highlight the increased difficulty of estimating MI in small-sample regimes and the consistent improvement in performance as the sample size grows.
Appendix D Self-Consistency Tests
D.1 Self-Consistency of Quantile Regression
[…]
  • Monotonicity: For a given fixed sample $(x,y)$, we sample a random uniform value of $\tau$ in $(0,1)$ and estimate $\hat{I}_{\tau}=\text{MIST}_{\text{QR}}(x,y,\tau)$. We repeat this process with $n$ values ($\tau_{1},\ldots,\tau_{n}$) and check whether $(\hat{I}_{\tau_{1}},\ldots,\hat{I}_{\tau_{n}})$ is an increasing sequence. We perform this operation on each data point of $\mathcal{M}_{\text{test}}$ using $n=1000$, and count the number of points for which monotonicity is not respected.
  • Bound checking: For a given fixed sample $(x,y)$ with true MI $I_{\text{true}}$, we compute $\hat{I}_{\text{lower}}=\text{MIST}_{\text{QR}}(x,y,\tau=0)$ and $\hat{I}_{\text{upper}}=\text{MIST}_{\text{QR}}(x,y,\tau=1)$. We perform this operation on each data point of $\mathcal{M}_{\text{test}}$, and count the number of points for which $I_{\text{true}}<\hat{I}_{\text{lower}}$ (lower bound failure) or $\hat{I}_{\text{lower}}<I_{\text{true}}$ (upper bound failure).
The results, given in Table 8, show that monotonicity breaks occur relatively rarely on both IMD and OoMD sets (1.1% and 1.4%, respectively). Further analysis reveals that monotonicity failures only occur on points with very low values of $I_{\text{true}}<0.05$. Bound failures happen at a moderate rate on the IMD dataset, and significantly more often for lower bounds (9.6%) than for upper ones (4.3%). When moving to OoMD distributions, however, those failure rates increase to 15.8% for lower bounds and 27.5% for upper ones, showing that this task remains difficult for our model.
IMD (1080 points) OoMD (1260 points)
Monotonicity breaks 15 (1.4%) 12 (1.2%)
Lower bound failures 104 (9.6%) 199 (15.8%)
Upper bound failures 46 (4.3%) 346 (27.5%)
[…]
D.2 Self-Consistency of the Learnt Quantity
[…]
  • The independence test checks that if $X\mathrel{\perp\!\!\!\perp}Y$, then $I(X;Y)=0$.
  • The data processing inequality checks that if $X,Y$ and $Z$ form a Markov chain $X\to Y\to Z$, then $I(X;Y)\geq I(X;Z)$ (further processing of the data cannot add new information).
  • The additivity over independent samples test checks that if $(X_{1},X_{2})\mathrel{\perp\!\!\!\perp}(Y_{1},Y_{2})$, then $I(X_{1},Y_{1};X_{1},Y_{2})=I(X_{1};X_{2})+I(Y_{1},Y_{2})$
(a) MSE Model
(b) QCQR Model, $\tau=0.95$
Figure 18: Independence Test: AA is an MNIST image, and $B_{r}$ contains the top $r$ rows of image AA. For a consistent estimator, the plotted MI ratio will be non-decreasing and approach 1 with an increasing number of shared rows.
(a) MSE Model
(b) QCQR Model, $\tau=0.95$
Figure 19: Data Processing Test: AA is an MNIST image, and $B_{r}$ contains the top $r$ rows of image AA. This test compares $I([A,A],[B_{r},B_{{}^{\prime}r-3^{\prime}}])$ to $I([A,B_{r}])$, where AA is an original MNIST image, $B_{r}$ is the first $r$ rows of $A$, and so $B_{{}^{\prime}r-3^{\prime}}$ is a further masked version of image AA. The test produces a ratio which should be near to 1 for all values of $r$, which amounts to checking that adding a further masked version of your image $B_{r}$ does not change the predicted MI. Images are encoded with PCA fit on 10K random samples, and error bars of $1\sigma$ reported, with 100 bootstrapped samples.
(a) MSE Model
(b) QCQR Model, $\tau=0.95$
Figure 20: Additivity Test: AA is an MNIST image, and $B_{r}$ contain the top $r$ rows of image AA. The premise of this test is that adding a second independent sample should double the mutual information estimate. The ratio compares $I([A_{1},A_{2}],[B_{1}r,B_{2}r])$ to $I([A,B_{r}])$, and should be near 2 for all values of $r$. Images are encoded with PCA fit on 10K random training samples. Error bars of $1\sigma$ are reported with 100 bootstrapped samples.

### v3

@@
 1 Introduction
-Mutual information (MI) is a measure of nonlinear statistical dependence between two random variables, quantifying how much knowing one variable reduces the uncertainty about the other (shannon1948mathematical; mackay2003information).
+Mutual information (MI) is a measure of nonlinear statistical dependence between two random variables, quantifying how much knowing one variable reduces the uncertainty about the other (Shannon, 1948; MacKay, 2003).
 For two continuous random variables $X$ and $Y$, the MI is written in terms of their joint distribution $P_{(X,Y)}$ and marginal distributions $P_{X}$ and $P_{Y}$:
 $\operatorname{I}(X;Y)=\int_{\mathcal{Y}}\int_{\mathcal{X}}P_{(X,Y)}(x,y)\log\left(\frac{P_{(X,Y)}(x,y)}{P_{X}(x)\,P_{Y}(y)}\right)\,dx\,dy.$
-Mutual information is one of the most widely applied information measures in data science, finding applications in feature selection (peng2005feature_selection; kwak2002_feature_selection), causality (butte1999_causality_genomics), representation learning (chen2016infogan; higgins2018towards_distentagled_representations; tishby2015_info_bottleneck; zhao2018information_autoencoders), reinforcement learning (oord2018_cpc; pathak2017curiosity), and generative modeling (alemi2016deep_variation_info_bottleneck; alemi2018gilbo; huang2020evaluating_lossy_compression).
+Mutual information is one of the most widely applied information measures in data science, finding applications in feature selection (Peng et al., 2005; Kwak and Choi, 2002), causality (Butte and Kohane, 1999), representation learning (Chen et al., 2016; Higgins et al., 2018; Tishby and Zaslavsky, 2015; Zhao et al., 2018), reinforcement learning (Oord et al., 2018; Pathak et al., 2017), and generative modeling (Alemi et al., 2016; Alemi and Fischer, 2018; Huang et al., 2020).
 Calculating MI requires access to the joint and marginal distributions, which are rarely known in practice.
 Consequently, most approaches infer MI from finite samples drawn from these distributions.
-Existing estimators fall into two main families.
-(1) Density estimators approximate the underlying densities, for example, using kernel density estimation, $k$-nearest neighbors, or variational autoencoders.
+Existing estimators fall into two main families: density estimators and density ratio estimators.
+Density estimators approximate the underlying densities, for example, using kernel density estimation, $k$-nearest neighbors, or variational autoencoders.
 MI can then be computed by substituting these estimates into one of MI’s definitions (such as Equation 1).
-Prominent density estimators include KSG (ksg) and MINDE (minde).
-(2) Density relationship estimators instead estimate the relationship between the joint and marginal distributions, often using variational lower bounds derived from the Donsker–Varadhan representation of the Kullback–Leibler divergence (see Appendix A.1 for these definitions).
-Notable examples include MINE (mine), NWJ (nwj), and SMILE (MINE_exp_variance_consistency_tests).
+Prominent density estimators include KSG (Kraskov et al., 2004) and MINDE (Franzese et al., 2023).
+Density ratio estimators instead estimate the relationship between the joint and marginal distributions, often using variational lower bounds.
+Notable examples include MINE (Belghazi et al., 2018), NWJ (Nguyen et al., 2010), and SMILE (Song and Ermon, 2019).
 In both families of approaches, neural methods have been developed to improve the estimator’s capacity to capture complex density functions or density relationships.
 Broadly, these neural methods have achieved better performance on moderately complex distributions and high-MI settings.
-However, both families of approaches struggle in challenging settings that are common in application, including high-dimension, limited samples, complex distributions, and high MI (bmi_benchmark; mcallester2020formal_limitations).
-Prominent examples of these settings include applications in genomics (klein2015droplet_genomics), neuroscience (ganguli2012compressed_neuroscience), astronomy (pandey2017much_astronomy), and machine learning (goldfeld2018estimating_info_flow).
+However, both families of approaches struggle in challenging settings that are common in practice: high dimensionality, limited samples, complex distributions, and high mutual information (Czyż et al., 2023; McAllester and Stratos, 2020).
+Examples of these settings include applications in genomics (Klein et al., 2015), neuroscience (Ganguli and Sompolinsky, 2012), astronomy (Pandey and Sarkar, 2017), and machine learning (Goldfeld et al., 2018).
 While previous benchmarking efforts have expanded the set of distribution families considered, MI estimators have rarely been empirically evaluated under these challenging conditions individually, and even more rarely in combination.
-Most empirical evaluations have studied estimation for MI below 1, for fewer than 5 distributions, with a minimum sample size above 10,000, and for higher dimensions limited to normal distributions.
-See Appendix A.3 for an overview of the empirical evaluation scale for existing methods.
+Most empirical evaluations have studied estimation for fewer than 5 distributions, with a minimum sample size of 10,000 or more, and in higher dimensions primarily for normal distributions (Appendix A.3).
+Existing work has instead primarily focused on theoretical guarantees of low bias, consistency with increasing samples, and the tightness of variational bounds (Belghazi et al., 2018; Nguyen et al., 2010; Song and Ermon, 2019).
 In this work, we propose a fully data-driven, empirical approach to designing mutual information estimators.
@@
 Our proposed model, MIST (Mutual Information estimation via Supervised Training), handles arbitrary sample sizes and dimensions, and features built-in uncertainty estimates.
-This meta-learning approach takes inspiration from amortized inference procedures such as simulation-based inference (SBI; pmlr-v89-papamakarios19a; cranmer2020frontier), amortized Bayesian inference (ABI; gonccalves2020training; elsemueller2024sensitivity; radev2023bayesflow; avecilla2022neural; pmlr-v235-gloeckler24a), and meta-statistical learning (peyrard2025meta).
-Our work focuses on the setting for MI estimators that was underexplored in previous work.
-Empirically, we conduct a large-scale evaluation that systematically explores the performance of our estimator in the most difficult regimes: with fewer samples (10-500), in higher dimensions (2-32), and across a wider range of MI values (0-40).
-We find that our learned estimator outperforms existing methods across these conditions, especially in low-sample regimes.
-We provide an empirical study of the behavior of estimators under increasing dimensionality, finding that our models require roughly half as many samples on average to provide reliable MI estimates as the best baseline.
+This meta-learning approach takes inspiration from amortized inference procedures such as simulation-based inference (SBI; Papamakarios et al., 2019; Cranmer et al., 2020), amortized Bayesian inference (ABI; Gonçalves et al., 2020; Elsemüller et al., 2024; Radev et al., 2023; Avecilla et al., 2022; Gloeckler et al., 2024), and meta-statistical learning (Peyrard and Cho, 2025).
+We focus on empirical settings underexplored in previous work but common in practice: fewer samples (10-500), higher dimensions (2-32), and a wider range of MI values (0-40) across 23 distribution types.
+We find that MIST substantially outperforms existing methods across these conditions, especially in low-sample regimes.
+We provide a detailed study of sample requirements under increasing dimensionality, finding that MIST requires approximately half as many samples on average to provide reliable MI estimates compared to the next best estimator.
 The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives, while inference is orders of magnitude faster than existing neural methods.
-Finally, the learned estimators generalize to unseen distributions and sample sizes.
+Finally, MIST generalizes effectively to unseen distributions and sample sizes.
 Beyond empirical performance, this meta-statistical framework establishes a new paradigm for constructing information-theoretic estimators.
 Because the estimator is fully differentiable and trainable, it can be embedded in larger learning pipelines, fine-tuned to specific data modalities, and systematically improved through curated meta-datasets.
-Leveraging MI’s invariance under invertible transformations, synthetic training distributions can, in the future, be adapted to arbitrary domains and modalities via normalizing flows (papamakarios2021normalizing_flows), paving the way for general-purpose, learned estimators of mutual information.
+Leveraging MI’s invariance under invertible transformations, synthetic training distributions can, in the future, be adapted to arbitrary domains and modalities via normalizing flows (Papamakarios et al., 2021), paving the way for general-purpose, learned estimators of mutual information.
 We provide an open-source library [^1] https://github.com/grgera/mist for training and evaluating meta-learned MI estimators, and include a detailed discussion of these opportunities and future work in Section 5.
@@
 Although these approaches vary in their mathematical foundations and computational strategies, they can be grouped into two main families: density estimators and density ratio estimators.
+See Appendix A.2 for a detailed discussion of previous estimators and our choices in categorizing them.
 Density Estimators.
-Density estimators compute mutual information estimates by performing intermediate estimates of density functions.
-The most straightforward approach is to use MI’s analytical definition, which requires estimates of the joint and marginal density functions.
-Histogram MI estimators approximate these density functions by partitioning the support of each variable into a set of bins and counting the frequencies in each (histogram_binning).
-Histogram estimators suffer from very poor scaling and are highly sensitive to the choice of bins, leading to research on optimizing the bin selection (adaptive_binning).
-Kernel density estimators (KDE; kde; textbook_kde) were later developed to help avoid this sensitivity by replacing fixed bins with kernel functions centered at each sample point.
-original_kl_density_estimator designed a method for entropy estimation using neighborhood distances, an approach which was later refined with weighted variants (weighted_kl_density_estimator).
-The Kraskov-Stögbauer-Grassberger estimator (KSG; ksg) extended this k-NN framework to mutual information estimation using MI’s expression in terms of conditional entropy (see Appendix A.1).
+Density estimators compute mutual information estimates by performing intermediate approximations of density functions.
+Histogram MI estimators approximate the joint and marginal density functions in 1 by partitioning the support of each variable into a set of bins and counting the frequencies in each (Fraser and Swinney, 1986).
+Histogram estimators suffer from very poor scaling and are highly sensitive to the choice of bins, leading to research on optimizing the bin selection (Darbellay and Vajda, 1999).
+Kernel density estimators (KDE; Moon et al., 1995; Silverman, 2018) were later developed to help avoid this sensitivity by replacing fixed bins with kernel functions centered at each sample point.
+Kozachenko (1987) designed a method for entropy estimation using neighborhood distances, an approach which was later refined with weighted variants (Berrett et al., 2019).
+The Kraskov-Stögbauer-Grassberger estimator (KSG; Kraskov et al., 2004) extended this k-nearest neighbor framework to mutual information estimation using MI’s expression in terms of conditional entropy (see Appendix A.1).
 KSG is one of the most broadly used MI estimators, especially for low-sample settings, and serves as a primary benchmark comparison in this work.
 Several approaches have explored the use of deep generative models for density estimation.
-One line of work uses normalizing flows to reconstruct density functions directly.
-However, these methods suffer from poor estimates even for data with simple structure (MINE_exp_variance_consistency_tests; dine_gaussian).
-The Barber-Agakov bound (BA; barber2004algorithm_BA_Bound) provides a variational lower bound on MI by introducing a variational approximation $q(x|y)$ to the true conditional $p(x|y)$.
-Several works have applied this bound using conditional generative models with different choices for the marginal distribution (alemi2016deep_variation_info_bottleneck; chalk2016relevant_BA_bound_application; kolchinsky2019nonlinear).
-Later work used VAEs to learn the joint and marginal density functions directly (MINE_exp_variance_consistency_tests).
-Most recently, MINDE (minde) trains conditional or joint diffusion models to learn score functions, then applies the Girsanov theorem to compute KL-divergence as the expected difference between joint and marginal scores.
-VCE (chen2025neural_vector_copulas) is a concurrent work to ours, and writes MI as the negative differential entropy of the vector copula density.
-Their approach uses flow models to estimate the marginal densities, then estimates the vector copula density using maximum likelihood estimation.
+One line of work uses VAEs or normalizing flows to reconstruct the joint and marginal density functions directly, an approach which suffers from high sample requirements and a limited ability to capture complex structure (Song and Ermon, 2019; Duong and Nguyen, 2023).
+The Barber-Agakov bound (BA; Barber and Agakov, 2004) provides a variational lower bound on MI by introducing a variational approximation $q(x|y)$ to the true conditional $p(x|y)$.
+Several works have applied this bound using conditional generative models with different choices for the marginal distribution (Alemi et al., 2016; Chalk et al., 2016; Kolchinsky et al., 2019).
+More recently, MINDE (Franzese et al., 2023) trains conditional or joint diffusion models to learn score functions, then applies the Girsanov theorem to compute KL-divergence as the expected difference between joint and marginal scores.
+VCE (Chen et al., 2025) is a concurrent work that expresses MI as the negative differential entropy of the vector copula density, which they estimate by first using normalizing flows to model the marginal densities, then applying maximum likelihood estimation to the copula.
 Another line of density estimators has focused on building MI estimators for complex settings by learning methods to simplify the distributions before applying classical estimators or computing MI directly.
-MIENF (mi_through_normalizing_flows) learns a pair of normalizing flows that apply MI-preserving transformations to the joint distribution, enabling closed-form MI calculation.
-Latent-MI (LMI; lmi_estimator) trains a network to compress data with minimal mutual information loss, then applies KSG to the low-dimensional embeddings.
+MIENF (Butakov et al., 2024) learns a pair of normalizing flows that apply MI-preserving transformations to the joint distribution, enabling closed-form MI calculation.
+Latent-MI (LMI; Gowri et al., 2024) trains a network to compress data with minimal mutual information loss, then applies KSG to the low-dimensional embeddings.
 These methods have improved abilities to handle high-dimensional data, but have large sample and computational requirements, and rely on the existence of simple or low-dimensional structure in high-dimensional data.
@@
 Density ratio estimators compute mutual information by estimating the relationship between the joint density $P(X,Y)$ and the product of marginal densities $P_{X}P_{Y}$.
-Since mutual information can be equivalently expressed as the Kullback-Leibler divergence $D_{KL}(P(X,Y)\|P_{X}P_{Y})$ (donsker_varadhan_kl_div_mi_definition), many density ratio methods leverage variational bounds on KL divergence.
+Since mutual information can be equivalently expressed as the Kullback-Leibler divergence $D_{KL}(P(X,Y)\|P_{X}P_{Y})$ (Donsker and Varadhan, 1975), many density ratio methods leverage variational bounds on KL divergence.
 Several methods have been developed by defining discriminative tasks that implicitly predict this density ratio.
-MINE (mine) trains a neural network critic to differentiate joint and marginal samples, optimizing the critic to maximize the Donsker-Varadhan (DV) lower bound.
-The Nguyen-Wainwright-Jordan estimator (NWJ; nwj) trains a critic model similarly, but uses an alternative variational bound which offers lower variance at the cost of a looser bound.
-InfoNCE (oord2018_cpc) trains a classifier critic to identify true joint samples among negative samples drawn from the marginals.
-SMILE (MINE_exp_variance_consistency_tests) combines the InfoNCE and NWJ bounds through a clip operation, interpolating between their respective bias-variance profiles to achieve more flexible trade-offs.
-The limitations of variational estimators are well-studied in mcallester2020formal_limitations, MINE_exp_variance_consistency_tests, and poole_variation_bounds_survey, and include high variance estimates, high sample requirements, and poor estimates for high MI settings.
-We refer the reader to poole_variation_bounds_survey for a detailed review of MI variational bounds.
-Evaluation of MI Estimators.
-Evaluating MI estimators is itself challenging: it requires distributions with known ground-truth MI and large sample sizes (required by most methods).
-Existing work has therefore primarily focused on theoretical guarantees of low bias, consistency with increasing samples, and the tightness of variational bounds (mine; nwj; MINE_exp_variance_consistency_tests).
-Empirical evaluations have been limited in scope.
-Early methods, such as binning and kernel density estimators, were tested on simple time-series data like sine waves (kde).
-Later neural methods began evaluation on a smaller set of synthetic distributions: MINE and SMILE were evaluated on 3–4 distributions, including high-dimensional Gaussians up to 20D (mine; MINE_exp_variance_consistency_tests).
-The majority of methods have been evaluated exclusively on MI values below 1, often on Gaussian or near-Gaussian distributions (see Table 3).
-Methods that consider higher dimensionalities do so largely by scaling low-dimensional distributions (assuming low-dimensional latent structure), or test only on high-dimensional Gaussians (lmi_estimator; minde).
-Recently, the BMI benchmark (bmi_benchmark) applied 20 invertible transformations to normal and Student’s t-distributions to create a broader set of distribution families with unique challenges, such as sparsity and long tails.
-This evaluation on more complex distributions highlighted the limitations of many existing estimators, leading authors to argue that multivariate normal distributions provided a overly optimistic view of estimator performance.
-BMI has been used by some of the most recent estimators, such as MINDE.
-We provide a description of our datasets in Appendix B, and an overview of existing evaluation scopes in Appendix A.3.
+MINE (Belghazi et al., 2018) trains a neural network critic to differentiate joint and marginal samples, optimizing the critic to maximize the Donsker-Varadhan (DV) lower bound.
+The Nguyen-Wainwright-Jordan estimator (NWJ; Nguyen et al., 2010) trains a critic model similarly, but uses an alternative variational bound which offers lower variance at the cost of a looser bound.
+InfoNCE (Oord et al., 2018) trains a critic to identify true joint samples among negative samples drawn from the marginals.
+SMILE (Song and Ermon, 2019) combines the InfoNCE and NWJ bounds through a clip operation, interpolating between their respective bias-variance profiles to achieve more flexible trade-offs.
+The limitations of variational estimators are well-studied in McAllester and Stratos (2020), Song and Ermon (2019), and Poole et al.
+(2019), and include high variance estimates, high sample requirements, and poor estimates for high MI settings.
+We refer the reader to Poole et al.
+(2019) for a detailed review of MI variational bounds.
+FMMI (Butakov et al., 2025) learns a coupling transform to map between the joint distribution and marginal product, using the expected divergence of the velocity field as the MI estimate.
+InfoNet (Hu et al., 2024) trains a neural network to predict the output of a trained MINE critic, removing the need to train a critic on each new set of samples.
+Like our work, this approach achieves substantial test-time efficiency gains through large-scale pretraining.
+However, our works differ in the learning objective: InfoNet learns to emulate an algorithm (MINE), whereas MIST learns an algorithm that allows it to predict the MI directly from samples.
 Machine Learning for Statistical Inference.
 Our work aligns with the broader research direction on learning inference procedures.
-A relevant related direction concerns neural processes (garnelo2018neural; garnelo2018conditional; kim2019attentive; Gordon:2020; markou2022practical; huang2023practical; BruinsmaMRFAVBH23), which can predict latent variables of interest from datasets (chang2025amortized) by leveraging transformers (pmlr-v162-nguyen22b) and Deep Sets (NIPS2017_f22e4747) to enforce permutation invariance (JMLR:v21:19-322).
-A related approach, known as prior-fitted networks, has demonstrated that transformers can be effectively repurposed for Bayesian inference (muller2022transformers) and optimization tasks (pmlr-v202-muller23a).
-Additionally, there is growing interest in using trained models to assist in statistical inference tasks (angelopoulos2023predictionpoweredinference) and optimization (NIPS2017_addfa9b7; NEURIPS2020_f52db9f7; NEURIPS2021_56c3b2c6; amos2023tutorialamortizedoptimization) like simulation-based inference (SBI; pmlr-v89-papamakarios19a; cranmer2020frontier), amortized Bayesian inference (ABI; gonccalves2020training; elsemueller2024sensitivity; radev2023bayesflow; avecilla2022neural; pmlr-v235-gloeckler24a) and the learning of frequentist statistical estimators (peyrard2025meta).
+A relevant related direction concerns neural processes (Garnelo et al., 2018b; a; Kim et al., 2019; Gordon et al., 2020; Markou et al., 2022; Huang et al., 2023; Bruinsma et al., 2023), which can predict latent variables of interest from datasets (Chang et al., 2025) by leveraging transformers (Nguyen and Grover, 2022) and Deep Sets (Zaheer et al., 2017) to enforce permutation invariance (Bloem-Reddy and Teh, 2020).
+A related approach, known as prior-fitted networks, has demonstrated that transformers can be effectively repurposed for Bayesian inference (Müller et al., 2022) and optimization tasks (Müller et al., 2023).
+Additionally, there is growing interest in using trained models to assist in statistical inference tasks (Angelopoulos et al., 2023) and optimization (Lueckmann et al., 2017; Liu et al., 2020; Simpson et al., 2021; Amos, 2023) like simulation-based inference (SBI; Papamakarios et al., 2019; Cranmer et al., 2020), amortized Bayesian inference (ABI; Gonçalves et al., 2020; Elsemüller et al., 2024; Radev et al., 2023; Avecilla et al., 2022; Gloeckler et al., 2024) and the learning of frequentist statistical estimators (Peyrard and Cho, 2025).
 3 Learning Framework
@@
 The methods available today decompose MI estimation into intermediate tasks of density function or density ratio estimation.
-To estimate a density function (or ratio) beyond very simple settings, these methods require large amounts of samples as part of inference.
-Even with access to a high number of samples, density estimation itself is a notoriously difficult and unsolved problem, particularly for high-dimensional or complex distributions (density_estimation).
-As a result, existing MI estimators still fail in many realistic data settings.
+To estimate a density function (or ratio) beyond simple settings, existing methods require large amounts of samples as part of inference.
+Even with access to a high number of samples, density estimation itself is a notoriously difficult and unsolved problem, particularly for high-dimensional or complex distributions (Mészáros et al., 2024).
+As a result, the MI estimators available today still frequently fail in many realistic data settings.
 They scale poorly to high-dimensional data and complex distributions, are unstable for low-sample settings, and require expensive inference procedures.
@@
 Our work offers an alternative philosophy for designing mutual information estimators.
-Instead of building estimators that compute MI with a pre-specified algorithm (e.g., approximating density functions or their ratios), we propose to predict MI directly from data.
+Instead of building estimators that compute MI with a pre-specified algorithm, such as approximating density functions or their ratios, we propose to predict MI directly from data.
 We take a fully empirical perspective, asking: Can a model learn a generalizable algorithm for predicting mutual information directly from samples?
-To learn this algorithm, we parameterize a mutual information estimator as a neural network that we call MIST (Mutual Information estimation from Supervised Training).
+To learn this algorithm, we parameterize a mutual information estimator as a neural network that we call MIST (Mutual Information estimation via Supervised Training).
 MIST takes as input a set of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $P_{XY}$, and outputs an estimate of $I(X;Y)$.
 The model is trained end-to-end on a large meta-dataset of synthetic joint distributions with known ground-truth MI values.
-We have reflected our framework in Figure 1 and described the meta-dataset design, architecture, and training choices below.
+We provide a visualization of our framework in Figure 1 and describe the meta-dataset design, architecture, and training choices below.
 The Learning Task
@@
 Training our meta-learned estimators in a supervised manner requires distributions with known MI.
-We generate these distributions using BMI, a library designed to create complex benchmarks for MI estimation.
+We generate these distributions using BMI (Czyż et al., 2023), a library designed to create complex benchmarks for MI estimation.
 BMI exploits the invariance of mutual information to homeomorphisms, applying invertible transformations to simple base distributions to produce more complex and challenging distributions.
@@
 From these distributions, we construct a training meta-dataset $\mathcal{M}_{\text{train}}$ and two test meta-datasets differing only in scale: $\mathcal{M}_{\text{test}}$, and $\mathcal{M}_{\text{test-extended}}$.
-These test meta-datasets contain a mix of distribution families found in $\Gamma_{\text{train}}$ (subsequently referred to as in-meta-distribution or IMD) or unseen at training time (out-of-meta-distribution or OoMD).
-Since MI is a function of two random variables, each meta-datapoint in our meta-datasets consists of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $\gamma\in\Gamma$, along with the known mutual information $I_{\gamma}(X;Y)$ as the label.
-The value of $n$ varies between $10$ and $500$ across each of our train and test meta-datasets, thus focusing on low-sample regimes.
+Both scales of test meta-datasets contain an in-meta-distribution (IMD) subset, in which the distribution family was seen in $\Gamma_{\text{train}}$, and an out-of-meta-distribution (OoMD) subset.
+Each meta-datapoint in our meta-datasets consists of $n$ paired samples $\{(x_{i},y_{i})\}_{i=1}^{n}$ drawn from a joint distribution $\gamma\in\Gamma$, along with the known mutual information $I_{\gamma}(X;Y)$ as the label.
+The value of $n$ varies between $10$ and $500$ across each of our train and test meta-datasets.
 The resulting training meta-dataset $\mathcal{M}_{\text{train}}$ contains 16 base distribution families and 625,000 meta-datapoints.
@@
 First, the model must process datasets of varying size while remaining invariant to sample ordering.
-The Transformer architecture is naturally permutation-invariant and has been a common choice to ingest datasets as inputs (transformer_invariance).
-The SetTransformer (set_transformer) architecture introduced the ISAB block to reduce the quadratic cost of attention by performing attention on a fixed number of learned inducing points.
-Finally, the SetTransformer++ architecture (set_transformer_plus_plus) further improves upon the original SetTransformer by introducing set normalization and additional residual connections that enable more stable training with deeper models.
+The Transformer architecture is naturally permutation-invariant and has been a common choice to ingest datasets as inputs (Xu et al., 2024).
+The SetTransformer (Lee et al., 2019) architecture introduced the ISAB block to reduce the quadratic cost of attention by performing attention on a fixed number of learned inducing points.
+Finally, the SetTransformer++ architecture (Zhang et al., 2022) further improves upon the original SetTransformer by introducing set normalization and additional residual connections that enable more stable training with deeper models.
 We therefore use SetTransformer++ as the basis of our architecture.
@@
 This is a long-standing challenge for MI estimators, which typically underestimate MI and fail in high-MI settings.
-We explored several normalization and multi-task prediction strategies (e.g., predicting normalized conditional entropies) to address this problem when designing our model, and provide our results in Appendix C.1.
+We explored several normalization and multi-task prediction strategies to address this problem when designing our model, and provide our results in Appendix C.1.
 We found that the simplest solution of directly predicting MI produced the best estimators, provided the meta-dataset includes a sufficient range of MI labels.
 4 Experiments
-Empirically, we find that our learned estimators outperform existing methods across sample sizes and dataset dimensions, particularly in low-sample settings (4.1).
-We provide an empirical study of the behavior of estimators under increasing dimensionality, finding that on average, our models require roughly half the number of samples as the best baseline to provide reliable MI estimates (4.2).
-The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives (4.3), while inference is orders of magnitude faster than existing neural methods (4.4).
-Finally, the learned estimators generalize to unseen distributions and sample sizes (4.5), and benefits from the mix of dimensions and size of the training dataset (4.6).
-4.1 Comparison with Existing Estimators
+Our experiments demonstrate that MIST outperforms existing methods across sample sizes and dimensions, particularly in low-sample regimes (4.1).
+MIST exhibits lower bias, lower variance, and higher CI coverage with increasing dimensions and samples (4.2).
+On average, MIST requires half the samples of the best baseline for reliable estimates in higher dimensions (4.3).
+The quantile-based intervals are well-calibrated and more reliable than bootstrap alternatives (4.4), while being orders of magnitude faster than existing neural methods (4.6).
+MIST generalizes to unseen distributions and sample sizes (4.5) and benefits from training on diverse dimensions and dataset sizes (4.7).
+4.1 MIST Outperforms Existing Estimators, Even on Unseen Distributions
 We begin by benchmarking our estimators against existing methods on a diverse set of synthetic distributions.
@@
 We use the smaller test set of 2,340 meta-datapoints to evaluate all baselines.
+In this challenging regime of low sample sizes, varying dimensionality, and non-Gaussian distributions, recent estimators struggle to outperform the classical KSG baseline.
+This aligns with recent findings demonstrating that evaluation on more complex settings reveals weaknesses in modern estimators (Czyż et al., 2023; Lee and Rhee, 2024).
+Our learned estimators substantially outperform all baselines.
+Relative to KSG (the next strongest method), our models achieve approximately $10\times$ lower error on distributions seen during training (IMD), and $\sim 5\times$ lower error on unseen distributions (OoMD).
+A breakdown by distribution family is provided in Appendix C.3.
+For subsequent large-scale evaluation, we focus on KSG as the baseline, since it is both computationally feasible on the extended test set and remains the strongest baseline estimator in Table 1, particularly in low-sample regimes.
+Table 1: MIST estimators substantially outperform existing methods across sample size ($n$) and distributions, including for distributions unseen during training (OoMD, right).
+Shown are the averaged MSE loss values (mean ± stddev) of each estimator, grouped by the number of samples $n$ given to the estimator, and using 100 bootstrap samples for the standard deviation.
 In-Meta-Distribution (IMD) Out-of-Meta-Distribution (OoMD)
@@
 InfoNCE $1.4\mathrm{e}^{2}$ $\pm 2.5\mathrm{e}^{2}$ $3.5\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$ $3.2\mathrm{e}^{1}$ $\pm 1.1\mathrm{e}^{2}$ $3.2\mathrm{e}^{2}$ $\pm 2.1\mathrm{e}^{3}$ $4.2\mathrm{e}^{1}$ $\pm 9.7\mathrm{e}^{1}$ $5.0\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$
+InfoNet $3.2\mathrm{e}^{1}$ $\pm 9.3\mathrm{e}^{1}$ $3.5\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$ $3.7\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$ $3.9\mathrm{e}^{1}$ $\pm 8.2\mathrm{e}^{1}$ $4.2\mathrm{e}^{1}$ $\pm 9.9\mathrm{e}^{1}$ $5.5\mathrm{e}^{1}$ $\pm 1.2\mathrm{e}^{2}$
+FMMI $2.8\mathrm{e}^{1}$ $\pm 8.4\mathrm{e}^{1}$ $2.5\mathrm{e}^{1}$ $\pm 9.8\mathrm{e}^{1}$ $1.7\mathrm{e}^{1}$ $\pm 6.5\mathrm{e}^{1}$ $3.3\mathrm{e}^{1}$ $\pm 7.4\mathrm{e}^{1}$ $2.8\mathrm{e}^{1}$ $\pm 7.7\mathrm{e}^{1}$ $3.1\mathrm{e}^{1}$ $\pm 8.2\mathrm{e}^{1}$
+VCE $1.5\mathrm{e}^{1}$ $\pm 3.9\mathrm{e}^{1}*$ $1.4\mathrm{e}^{1}$ $\pm 4.3\mathrm{e}^{1}$ $1.6\mathrm{e}^{1}$ $\pm 3.6\mathrm{e}^{1}$ $2.6\mathrm{e}^{1}$ $\pm 6.5\mathrm{e}^{1}*$ $2.4\mathrm{e}^{1}$ $\pm 6.6\mathrm{e}^{1}$ $2.4\mathrm{e}^{1}$ $\pm 5.9\mathrm{e}^{1}$
 DV $3.0\mathrm{e}^{19}$ $\pm 2.1\mathrm{e}^{20}$ $1.1\mathrm{e}^{9}$ $\pm 2.2\mathrm{e}^{10}$ $8.5\mathrm{e}^{6}$ $\pm 6.7\mathrm{e}^{7}$ $1.2\mathrm{e}^{19}$ $\pm 6.3\mathrm{e}^{19}$ $8.7\mathrm{e}^{6}$ $\pm 5.1\mathrm{e}^{7}$ $6.1\mathrm{e}^{6}$ $\pm 5.8\mathrm{e}^{7}$
@@
 MISTQR $3.1\mathrm{e}^{0}$ ± $11.7$ $3.1\mathrm{e}^{-1}$ ± $1.5$ $1.2\mathrm{e}^{-1}$ ± $0.3$ $9.5\mathrm{e}^{0}$ ± $17.1$ $9.9\mathrm{e}^{0}$ ± $24.6$ $1.0\mathrm{e}^{1}$ ± $25.4$
-Table 1: Comparison of MI estimation baselines and our proposed methods across different distribution families (IMD and OoMD) and varying numbers of samples per meta-datapoint on the $\mathcal{M}_{\text{test}}$.
-Our estimators substantially outperform existing methods across sample size and distribution settings, including for distributions unseen during training.
-Shown are the averaged MSE loss values (mean ± stddev) of each estimator, grouped by the number of samples $n$ given to the estimator.
-The standard deviation is computed from 100 bootstrap resamples for each regime.
-In this challenging regime of low sample sizes, varying dimensionality, and non-Gaussian distributions, recent estimators struggle to outperform the classical KSG baseline.
-This aligns with recent findings demonstrating that evaluation on more complex settings reveals weaknesses in modern estimators (bmi_benchmark; lee2024a).
-Our learned estimators substantially outperform all baselines.
-Relative to KSG (the next strongest method), our models achieve approximately $10\times$ lower error on distributions seen during training (IMD), and $\sim 5\times$ lower error on unseen distributions (OoMD).
-A breakdown by distribution family is provided in Appendix C.3.
-For subsequent large-scale evaluation, we focus on KSG as the baseline, since it is both computationally feasible on the extended test set and remains the strongest baseline estimator in Table 1, particularly in low-sample regimes.
+* VCE failed to produce an estimate for 99/216 points (118/252) in the IMD (OoMD) subsets.
+4.2 MIST Exhibits Lower Bias and Variance Across Settings
+Figure 2: Predicted MI as a function of the true MI on $\mathcal{M}_{\text{test-extended}}$. For each estimator, predictions are aggregated into bins of true MI values.
 […]
-Figure 2: Heatmaps providing a detailed analysis of the performance of the two proposed MI estimators and the KSG baseline.
-The data dimensions are grouped into three categories, and within each group, meta-datapoints are sorted by the number of samples.
-Shown are the average values of MSE, bias, variance, and confidence interval (CI cov.) width across $\mathcal{M}_{\text{test-extended}}$.
-CI coverage denotes the fraction of samples per group for which the true MI lies within the 95% bootstrap confidence interval.
 […]
-Figure 3: Predicted MI as a function of the true MI on $\mathcal{M}_{\text{test-extended}}$. For each estimator, predictions are aggregated into bins of true MI values.
 One of the most well-documented failure modes of existing MI estimators is their negative bias, in which estimators substantially underestimate predictions for distributions with large MI.
@@
 While KSG’s estimates plateau quickly, our estimates closely follow the true MI values.
-4.2 Scaling Behavior
+Figure 3: MIST estimators widely outperform KSG, remaining substantially less biased and lower variance across dimensions groups and sample sizes in $\mathcal{M}_{\text{test-extended}}$.
+CI coverage denotes the fraction of samples per group for which the true MI lies within the 95% bootstrap confidence interval.
+4.3 MIST Requires Fewer Samples to Scale to Higher Dimensions
 Traditional evaluations of MI estimators have been constrained to relatively simple distributions and high-sample regimes.
-While recent efforts such as the BMI benchmark have enabled more comprehensive studies on complex distributions, systematic empirical analysis of scaling behavior has remained limited, largely due to the computational expense of evaluating existing methods.
+While recent efforts have enabled more comprehensive studies on complex distributions, systematic empirical analysis of scaling behavior has remained limited, largely due to the computational expense of evaluating existing methods.
 Leveraging our larger test set $\mathcal{M}_{\text{test-extended}}$, we conduct a detailed empirical study of estimator performance across a wide range of dimensions and sample sizes.
@@
 […]
-4.3 Calibration
+4.4 MIST Quantile-Based Intervals are More Reliable Than Bootstrapping
 […]
@@
 […]
-4.4 Inference Efficiency
+4.5 MIST Generalizes to Unseen Distributions, Dimensions, and Sample Sizes
+The previous sections demonstrate the unique ability of meta-learned estimators to generalize to unseen distribution families (OoMD) at test-time.
+The strength of these results motivates a bolder question: if meta-learned MI algorithms can generalize to unseen distributions, can they also generalize to dimensions and sample sizes beyond their training experience?
+We conduct a controlled study by training models on only the subset of $M_{\text{train}}$ with $D<16$ and $n<300$, then evaluating on the full test set $\mathcal{M}_{\text{test-extended}}$ which includes unseen higher dimensions ($D>16)$, unseen larger sample sizes ($n>300$), and unseen new distributions (OoMD).
+Figure 6 presents loss heatmaps for in-distribution (IMD) and out-of-distribution (OoMD) test sets.
+We find that our estimators generalize well to new dimensions and sample sizes for distributions that are in the training meta-distribution (IMD heatmap).
+Similarly, we find that MIST models generalize well to new distributions when the data dimensionality is similar to that during training (OoMD heatmap, left half).
+However, they struggle to generalize to unseen distributions when the data also has higher dimensionality than any seen during training (OoMD heatmap, right half).
 […]
+4.6 MIST Achieves Large Gains in Inference Efficiency
+Figure 7: Average inference time for each method with increasing sample size.
+[…]
+The combination of fast inference and efficient uncertainty quantification (as discussed in Section 4.4) makes our approach practical for large-scale applications where repeated MI estimation is required, such as hyperparameter tuning, feature selection, or iterative model training (see Appendix E for an example of use in feature selection).
 We estimated the inference time of each baseline on the entirety of $\mathcal{M}_{\text{test-extended}}$ and report the results in Appendix C.4.
 The high computational costs (>> 1 week) of evaluating all baselines except for CCA and KSG on $\mathcal{M}_{\text{test-extended}}$, along with the poor performance measured for CCA on the $\mathcal{M}_{\text{test}}$ for low sample sizes, further motivate our selection of KSG as the only baseline for other experiments in this work.
-Figure 6: Inference time for each method as a function of the sample size. Results are averaged over 100 inferences (5 for MINDE), each performed over 10 meta-datapoints.
-4.5 Generalization to Unseen Distributions, Dimensions, and Sample Sizes
-The previous sections demonstrate that the learned algorithm generalizes effectively to unseen distribution families, which is evaluated at dimensions and sample sizes seen during training.
-We find that by learning to infer MI across diverse distributions and sample regimes, our estimators learn appropriate regularization strategies that transfer even to distribution families not encountered during training.
-The strength of these results motivates a bolder question: if meta-learned MI algorithms can generalize to unseen distributions, can they also generalize to these distributions at dimensions and sample sizes beyond their training experience?
-We conduct a controlled study by training models on restricted subsets of the training data, specifically training on meta-datapoints in $M_{\text{train}}$ with a dimensionality $D<16$ and a sample size $S<300$.
-We evaluate performance on our full test meta-dataset $\mathcal{M}_{\text{test-extended}}$, which therefore includes distributions with higher dimensionalities and sample sizes and new distributions.
-Figure 7 presents two loss heatmaps, one for $\mathcal{M}_{\text{test-extended,IMD}}$ and one for $\mathcal{M}_{\text{test-extended,OoMD}}$.
-We find that our estimators generalize well to new dimensions and sample sizes for distributions that are in the training meta-distribution (IMD heatmap).
-Similarly, we find that our models generalize well to new distributions when the data dimensionality is similar to that during training (OoMD heatmap, left half).
-However, they struggle to generalize to unseen distributions when the data also has higher dimensionality than any seen during training (OoMD heatmap, right half).
+On our multidimensional test set, InfoNet’s slicing overhead offsets its efficiency gains over optimized methods like MINE.
+4.7 Training on Multiple Dimensions Benefits High-Dimensional Estimation
+In contrast to existing paradigms, MIST models learn over a diverse meta-distribution rather than fitting an individual dataset of interest.
+This approach opens many relevant empirical questions regarding the design of the meta-distribution.
+One of the most prominent design choices is the choice to train over many dimensions or to train a distribution-specific model.
+While training over many dimensions has the benefit of generalized use, practitioners may know the dimension of their data a priori.
 […]
-4.6 Impact of Variable Input Dimensions
-A natural question that arises from our approach is the impact of including multiple dimensions in model training.
-Of course, training on a mix of dimensions allows the development of a general-purpose estimator that can achieve high performance across settings, as shown in the above experiments.
-However, there may be settings where practitioners know their data dimension apriori, and may consider producing a training dataset of only that dimension.
-To study these questions, we trained separate (specialized) copies of MIST on subsets of $\mathcal{M}_{\text{train}}$ filtered to contain only meta-datapoints with a fixed input dimension (e.g., 4, 8, 16 or 32).
+To study these questions, we trained separate (specialized) MIST models on subsets of $\mathcal{M}_{\text{train}}$ filtered to contain only meta-datapoints with a fixed input dimension (e.g., 4, 8, 16 or 32).
 This filtering procedure reduces the number of training data points substantially (by a factor of 31).
-To provide a comparison with the same number of samples, we also trained an additional (reduced data) MIST model on a uniformly down-sampled version of $\mathcal{M}_{\text{train}}$ to match the training size of the individual dimension models.
+To provide a comparison with the same number of samples, we also trained additional (reduced data) MIST models on a uniformly down-sampled version of $\mathcal{M}_{\text{train}}$ to match the number of samples seen by the individual dimension models.
 By comparing the base MIST (full training set) with the specialized models (only one dimension), we can measure the impact of scaling both the dimensions and dataset size.
@@
 In Figure 8, we provide these comparisons for each input dimension in $\mathcal{M}_{\text{test-extended}}$.
-[…]
 […]
@@
 Our estimator’s behavior is learned, rather than designed, and therefore does not carry the theoretical consistency guarantees of classical MI estimators—namely, that the estimate converges to the true MI as $n\to\infty$.
-We studied our estimator’s behavior on self-consistency tests from MINE_exp_variance_consistency_tests in Appendix D, and find that while our estimators are approximately self-consistent under the independence test, they are not self-consistent and produce high variance prediction ratios for the data processing and additivity tests.
-We believe that a learning-based MI framework offers opportunity for designing models, datasets, and losses to induce such desired behavior characteristics, but we believe it is important to recognize that our estimators do not carry these properties innately.
+We studied our estimator’s behavior on self-consistency tests from Song and Ermon (2019) in Appendix D, and find that while our estimators are approximately self-consistent under the independence test, they are not self-consistent and produce high variance prediction ratios for the data processing and additivity tests.
+In Appendix E, we study the utility of MIST predictions downstream in a task of learned feature selection, using MINE as a baseline.
+In this setting, we find that MIST-guided learning was more accurate, lower-variance, and more stable across a range of sample sizes and noise levels.
+A learning-based MI framework offers opportunity for designing models, datasets, and losses to induce desired behavior characteristics, but we believe it is important to recognize that our estimators do not carry these properties innately and empirical evaluation is required.
+Another important consequence of MIST being a learned estimator is that the range of MI values predicted by MIST are limited to those seen during training.
 In high-sample settings, our learned estimator cannot directly leverage all available samples during one inference pass and other estimators may be preferred if time and compute resources are available.
@@
 These limitations parallel the general challenges faced by machine learning models in other domains, highlighting the importance of careful validation and uncertainty quantification.
-Controllability and adaptability.
-A key strength of this framework lies in its controllability.
-Our estimators’ behavior are largely determined by the synthetic datasets created; practitioners can customize, interrogate, and design datasets to improve estimator quality and reliability.
-Since mutual information is invariant under invertible transformations, our meta-dataset can be mapped to a target modality using normalizing flows without altering its ground-truth MI.
-This property enables the generation of diverse synthetic meta-datasets that reflect the characteristics of arbitrary data domains, such as images, text, tabular data, or multimodal signals.
-By tailoring the training data through such transformations, one can design estimators specialized for particular classes of distributions or downstream applications.
-This mechanism opens the door to what may be viewed as a foundational model for information-theoretic estimation, one that can be adapted via fine-tuning to specific tasks and modalities.
+[…]
 Computational efficiency.
@@
 Donsker-Varadhan
-A commonly used expression of MI is in terms of the Kullback-Leibler (KL) divergence between the joint distribution and the product of marginals (donsker_varadhan_kl_div_mi_definition):
+A commonly used expression of MI is in terms of the Kullback-Leibler (KL) divergence between the joint distribution and the product of marginals (Donsker and Varadhan, 1975):
 $I(X;Y)=D_{\text{KL}}(P_{(X,Y)}\|P_{X}\otimes P_{Y})$ (5)
@@
 A.2 MI Estimators
-We provide an overview of prominent MI estimators for reference in Table 2, categorized by their prediction category, the MI definition considered, and the estimation task.
-Previous works such as MINE_exp_variance_consistency_tests have described variational estimators as either discriminative or generative, with discriminative methods learning the ratio between joint and marginal densities directly and generative methods estimating the densities separately before computing the density ratio.
-We broaden this categorization into density and density ratio estimators.
-Method Category $I(X;Y)$ Estimation Task
+We provide an overview of prominent MI estimators for reference in Table 2, categorized by the type of estimation they perform, the MI definition considered, and the specific estimation task.
+Previous works have often adopted the terminology introduced by Song and Ermon (2019).
+This framework categorizes variational estimators as either discriminative or generative: generative methods estimate the densities individually and discriminative methods estimate the ratio between joint and marginal densities.
+Given the diversity of recent estimation approaches, we broaden this categorization into density and density ratio estimators.
+In this categorization, density estimators estimate MI by approximating the form or properties of densities $P_{X},P_{Y},\text{and }P_{X,Y}$.
+This includes learning the densities explicitly with a generative model (GM, KDE, VCE), approximating densities through k-nearest-neighbor distances (KSG), learning flows to transform the marginal distributions to a tractable distribution (MIENF), or learning score functions (MINDE).
+In contrast, density ratio estimators estimate MI by modeling the relationship between densities, primarily in the form of the density ratio $log(\frac{P_{X,Y}}{P_{X}P_{Y}})$, without modeling the densities individually.
+This estimation can be performed by learning classifiers (’critics’) to distinguish joint and marginal samples (MINE, SMILE, InfoNCE), or learning coupling transforms that map between the joint and marginal product (FMMI).
+We chose this naming convention to better capture the diversity of existing methods and highlight the core task performed by each method.
+We hope that this naming convention can help avoid reader confusion between the modeling framework and the core MI estimation task.
+For example, works like FMMI are generative in terms of modeling choices (flows), but discriminative in terms of the MI estimation task.
+Similarly, many classical estimators are not generative in terms of modeling choices (binning, KSG), but are generative in terms of MI estimation.
+To improve clarity, we group methods by their estimation task (estimate a density, or a density ratio), regardless of the approach used to perform the estimation.
+Still, this categorization, as with any abstraction, removes some nuance.
+We provide this appendix section as a resource for those looking to gain a better holistic understanding of existing approaches, and encourage readers to read the respective papers for precise details.
+Table 2: Summary of prominent methods for estimating mutual information, along with the MI definition used and their estimation tasks.
+Method Type $I(X;Y)$ Estimation Task
 Binning Density $\sum_{x,y}p(x,y)\log\frac{p(x,y)}{p(x)p(y)}$ Estimate $p(x,y)$, $p(x)$, $p(y)$ via histograms
-KSG Density =H(X)+H(Y)H(X,Y)=H(X)+H(Y)-H(X,Y) Implicitly estimate H()H(*) by writing H()H(*) in terms of KNN distances
+KSG Density =H(X)+H(Y)H(X,Y)=H(X)+H(Y)-H(X,Y) Implicitly estimate $H(\cdot)$ by writing MI’s entropy definition in terms of KNN distances
 KDE Density $\sum_{x,y}p(x,y)\log\frac{p(x,y)}{p(x)p(y)}$ Estimate $p(x,y)$, $p(x)$, $p(y)$ via kernels
-MIENF Density $I(f_{X}(X);f_{Y}(Y))$ Learn flows $f_{X}$, $f_{Y}$ to Gaussian marginals
-LMI Density $\approx I(Z_{x};Z_{y})$ Learn neural networks for embeddings $Z_{x}=g_{X}(X)$, $Z_{y}=g_{Y}(Y)$, then apply KSG
+MIENF Density $\approx I(f_{X}(X);f_{Y}(Y))$ Learn flows $f_{X}$, $f_{Y}$ to Gaussian marginals
+LMI Density $\approx I_{KSG}(Z_{x};Z_{y})$ Learn embeddings $Z_{x}=g_{X}(X)$, $Z_{y}=g_{Y}(Y)$ via cross-predictive autoencoders $g_{X},g_{Y}$, then apply KSG
 GM Density $\int\int p(x,y)\log\frac{p(x,y)}{p(x)p(y)}dx\,dy$ Estimate $p(x,y)$, $p(x)$, $p(y)$ via VAEs (or flows)
-MINDE Density $D_{KL}(p(x,y)\|p(x)p(y))$ Learn scores $\nabla\log p(x,y)$, $\nabla\log p(x)$ $\nabla\log p(y)$
+MINDE Density $\begin{aligned} &=D_{KL}(p(x,y)\|p(x)p(y))\\ &\approx\|\nabla\log p(x,y)-\nabla\log p(x)p(y)\|^{2}\end{aligned}$ Learn scores $\nabla\log p(x,y)$, $\nabla\log p(x)$, $\nabla\log p(y)$ with a diffusion model
 VCE Density $H[c(\mathbf{u}_{X},\mathbf{u}_{Y})]$ Learn marginal ranks $\mathbf{u}_{X}$, $\mathbf{u}_{Y}$ with flows, estimate vector copula $c$ with MLE over mixture of Gaussian copulas
-MINE Density ratio $\geq\sup_{\theta}\mathbb{E}_{P_{XY}}[T_{\theta}]-\log(\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}}])$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approx.
-$\log(\frac{p(x,y)}{p(x)p(y)})$
-NWJ Density ratio $\geq\sup_{\theta}\mathbb{E}_{P_{XY}}[T_{\theta}]-\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}-1}]$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approx.
-$\log\frac{p(x,y)}{p(x)p(y)}$
-InfoNCE Density ratio $\geq\mathbb{E}_{(x,y)\sim P_{n}}\left[\log\frac{f_{\theta}(x,y)}{\frac{1}{n}\sum_{j=1}^{n}f_{\theta}(x,y_{j})}\right]$ Learn classifier critic $f_{\theta}(x,y)\propto\frac{p(x,y)}{p(x)p(y)}$
-SMILE Density ratio $\geq\mathbb{E}_{P}[T_{\theta}]-\log\mathbb{E}_{Q}[\text{clip}{\tau}(e^{T_{\theta}})]$ * Learn classifier critic $f_{\theta}(x,y)$ to maximize bound.
-MIST From Samples $\approx F(x,y)$ Learn estimator $F$ by direct optimization
-Table 2: Summary of prominent methods for estimating mutual information, along with the MI definition used and their estimation tasks.
-$\text{clip}_{\tau}(z)\in[e^{-\tau},e^{\tau}]$.
+MINE Density Ratio $\geq\sup_{\theta}\mathbb{E}_{P_{X,Y}}[T_{\theta}]-\log(\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}}])$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approximating $\log(\frac{p(x,y)}{p(x)p(y)})$
+NWJ Density Ratio $\geq\sup_{\theta}\mathbb{E}_{P_{X,Y}}[T_{\theta}]-\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{T_{\theta}-1}]$ Learn critic $T_{\theta}(x,y)$ to maximize bound, approximating $\log(\frac{p(x,y)}{p(x)p(y)})+1$
+InfoNCE Density Ratio $\geq\mathbb{E}_{(x,y)\sim P_{X,Y}}\left[\log\frac{f_{\theta}(x,y)}{\frac{1}{n}\sum_{j=1}^{n}f_{\theta}(x,y_{j})}\right]$ Learn classifier critic $f_{\theta}(x,y)\propto\frac{p(x,y)}{p(x)p(y)}$
+SMILE Density Ratio $\geq\mathbb{E}_{P_{X,Y}}[T_{\theta}]-\log\mathbb{E}_{P_{X}\otimes P_{Y}}[\text{clip}_{\tau}(e^{T_{\theta}})]$ * Learn classifier critic $T_{\theta}(x,y)$ to maximize bound, approximating $\log(\frac{p(x,y)}{p(x)p(y)})$
+FMMI** Density Ratio $\begin{aligned} &=-\mathbb{E}_{T\sim\mathcal{U}[0,1]}[\text{div }v(Z_{T},T)]\\ &Z_{0}\sim P_{X}\otimes P_{Y},\,Z_{1}\sim P_{X,Y}\end{aligned}$ Train velocity network $v(z,t)$ via flow matching, estimate MI as negative expected divergence of $v$
+InfoNet Density Ratio $\geq\sup_{\theta}\mathbb{E}_{P_{X,Y}}[\theta]-\log(\mathbb{E}_{P_{X}\otimes P_{Y}}[e^{\theta}])$ Pretrain network $F(x,y)\rightarrow\theta$, where $\theta$ is MINE’s optimal discriminant as as a look-up table
+MIST Meta- Learned $\approx F(x,y)$ Learn estimator $F$ by direct optimization
+* $\text{clip}_{\tau}(z)\in[e^{-\tau},e^{\tau}]$.
+** FMMI learns a coupling transform between joint and marginal distributions.
+This estimation implicitly encodes the density ratio, which could be derived from the velocity field with an integration.
+However, FMMI does not perform this integral and does not estimate the density ratio at each point.
+Instead, FMMI estimates MI as an expectation of the field’s divergence.
+FMMI authors describe their method as a hybrid method.
+We categorize FMMI as a density ratio method due to the focus on learning the relationship between joint and marginal distributions, rather than modeling the densities individually.
 A.3 Empirical Evaluation Scope of MI Estimators
 In Table 3, we provide a summary of how prominent estimators were empirically evaluated.
-We focus our summary on evaluations of the estimator directly, rather than the performance of an estimator in a downstream tasks.
-Similarly, we only include the evaluations of estimators in the papers introducing the method, as we are most interested in how estimator quality is determined.
+We only include the evaluations of estimators in the papers introducing the method, as we are most interested in how estimator quality is determined.
+This summary focuses on the experiments evaluating the estimator directly, and excludes evaluations in downstream settings, where ground truth MI values are not available.
+Table 3: Summary of the empirical evaluation scope of prominent mutual information estimators.
+Comma-separated entries indicate values for different experimental settings (e.g., KSG has four experiments).
+Dashes indicate data that is not available or not specified.
 Method Data Type # $\mathcal{D}$ # $\not\sim\mathcal{N}$ $d_{\max}$ $I_{\max}$ $n_{\min}$
 Binning Time-Series 1 1 1 - 65K
-KDE Time-Series 4 4 1 - 400, 2K, 4K
-KSG Synthetic 4 3 3, 8 0.83,1,1.2 125, 10K, 50K
-MIENF Synthetic 6 5 32, 1024 10 10K
-LMI Synthetic,MNIST,PLM 3 2 5K,784,1024 1 2K,5K
+KSG Synthetic 1,3,1,1 0,3,0,0 2,1,3,8 0.83,1,1,1.2 125,1K,10K,50K
+KDE Time-Series 4 4 1 - 400
+MIENF Synthetic 1,5 0,5 1024,32 10,10 10K,10K
+LMI Synthetic,Images,PLM 1,1,1,1 0,0,1,1 5K,50,784,1024 2,1,1,1 2K,100,5K,4.4K
 GM Synthetic 2 1 20 10 256K
-MINDE Synthetic 20, 3 17, 2 25, 9 1, 5 100K
-VCE Synthetic,Images,Text 7,2,2 6,1,2 64,16,16 16,7,2.1 10K,10K,4K
-MINE Synthetic 2 0 2, 20 2, 40 -
+MINDE Synthetic 19,3 16,2 25,9 1,5 100K,-
+VCE Synthetic,Images,Text 5,5,2,2 4,3,1,2 64,64,16,16 16,14,7,2.1 10K,10K,10K,4K
+MINE Synthetic 2,3 0,3 20,2 40,- -,-
 NWJ - - - - - -
@@
 SMILE Synthetic 2 1 20 10 256K
+FMMI Synthetic 4 2 160 115 10K
+InfoNet Synthetic 1,20,10,3 0,19,9,0 1,1,1,1281 $0.9,$-,- 2K,2K,2K,50
 MIST Synthetic 23 19 32 40 10
-Table 3: Summary of the empirical evaluation scope for prominent mutual information estimators.
-Comma-separated entries indicate values for different experimental settings (e.g., KDE has three experiments with minimum sample size considered 400, 2K, and 4K, respectively).
-Dashes indicate data that is not available or not specified.
+1 Using slicing, which takes an expectation over random 1D projections
 Early methods for mutual information estimation, such as binning, were evaluated primarily using simulated time-series data.
 These commonly included sine and cosine functions, as well as chaotic dynamical systems such as the Lorenz and Rossler attractors (where the task is phase reconstruction, and no ground truth is available).
-In the initial histogram binning work (histogram_binning), the Rossler attractor was used.
+In the initial histogram binning work (Fraser and Swinney, 1986), the Rossler attractor was used.
 For this work, we include it in the table as having one 1D distribution, which is non-Gaussian, with no ground truth MI reported (and therefore no maximum).
 The experiment calculated the mutual information over 65,536 points.
-KDE (kde) was similarly evaluated with time-series data.
+KDE (Moon et al., 1995) was similarly evaluated with time-series data.
 KDE was evaluated on 1D synthetic sine waves, an autoregressive function of a sine and a Gaussian distribution, the Lorenz attractor, and the Rossler attractor.
@@
 Estimators are evaluated on the quality of the reconstruction, with no ground truth MI available.
-We include KDE in the table as having 4 distributions (sine, AR, Lorenz, Rossler), 4 of which are non-Gaussian, and all 1D.
-No ground truth MI values are present, and the samples included 400, 2K, 4K.
-KSG (ksg) performed evaluation on 2D correlated Gaussians with unit variance and zero mean, testing sample sizes of 125, 250, 500, 1K, 2K, 4K, 10K, and 20K.
+KSG (Kraskov et al., 2004) performed evaluation on 2D correlated Gaussians with unit variance and zero mean, testing sample sizes of 125, 250, 500, 1K, 2K, 4K, 10K, and 20K.
 The MI was 0.830366.
@@
 The minimum sample size in these experiments was 125 for the 1-D Gaussian, 10K for other distributions, and 50K for larger dimensions.
-MIENF (mi_through_normalizing_flows) first evaluated their estimator on 2D correlated Gaussian distributions mapped into higher-dimensional images (16x16 and 32x32), evaluated on 10K samples, and with MI values between 0 and 10.
+MIENF (Butakov et al., 2024) first evaluated their estimator on 2D correlated Gaussian distributions mapped into higher-dimensional images (16x16 and 32x32), evaluated on 10K samples, and with MI values between 0 and 10.
 In their second experiment, authors evaluated their method on five non-compressible distributions (uniform, smoothed uniform, student (dof = 4), arcsignh (student) student with dof = 1,2).
@@
 The MI was set to 10 for all distributions and dimensions.
-Latent-MI (LMI; lmi_estimator) was evaluated on multivariate Gaussian distributions under varying original data dimensionality and latent dimensionality learned before applying KSG.
+Latent-MI (LMI; Gowri et al., 2024) was first evaluated on multivariate Gaussian distributions under varying original data dimensionality and latent dimensionality (the dimensionality of the KSG input).
 Their multivariate Gaussian experiment considered up to 10 latent dimensions and 5K original dimensions, with 2K samples.
+The MI ranged from 0 to 2 bits.
 In a second experiment, the authors studied the sample requirements of their estimators.
 For this experiment, authors used multivariate Gaussians from dimension 1 to 50, each with one correlated dimension, and an MI of 1.
-Authors varied the sample size up to 4K.
+Authors varied the sample size from 100 to 10K.
 In a third experiment, MNIST and protein embeddings were resampled to align with a Bernoulli vector with set correlation.
-For MNIST, the dimension was 784, and the sample size was 5K.
-For the protein embeddings, the dimension was 1024 and the sample size was 4402.
-In both cases, datasets were generated with MI values between 0 and 1.
-For GM and SMILE (MINE_exp_variance_consistency_tests), evaluation was performed on two tasks.
-In the first task, authors evaluated their methods with 20-D Gaussians with correlation $p$, and in the second task, the authors applied a cubic operation on these Gaussians.
-This set-up was introduced in (mine).
-MI values were increased from 0 to 10 at step intervals during the critic training process.
+We consider the MIST and protein embedding experiments separately.
+For MNIST, the dimension was 784, and the sample size was 5K, and the dataset was generated with MI values between 0 and 1.
+For the protein embeddings, the dimension was 1024, the sample size was 4402, and the dataset was generated with MI values between 0 and 1.
+For GM and SMILE (Song and Ermon, 2019), evaluation was initially performed two distributions.
+The first distribution was a 20-D Gaussian with correlation $p$ (borrowing from a setup in (Belghazi et al., 2018)), and the second was this distribution with a cubic operation applied.
+Authors follow a setup from (Poole et al., 2019), in which MI values were increased from 0 to 10 at step intervals during the critic training process.
 Training occurred over 20K steps, with MI increasing by 2 every 4K steps.
@@
 Since the estimator saw 4K steps of batch size 64 for each MI (2, 4, 6, 8, 10), we list the minimum samples seen as 4,000 x 64 = 256K.
-This work also includes consistency tests constructed by adapting MNIST images, but this setting does not carry ground-truth MI values (only the ideal ratio of estimates).
-MINDE (minde) used the BMI benchmark (bmi_benchmark) at larger sample sizes (100K rather than 10K).
-For our table, we consider the number of unique distribution types considered; for BMI, there are 20 unique distribution tasks after excluding tasks that are higher-dimensional versions of another.
-The highest MI value used was 5, and the maximum dimension considered was 25.
-MINDE also performs image consistency experiments as in SMILE, but this setting does not have a ground truth value.
-VCE chen2025neural_vector_copulas evaluate estimators on five synthetic distributions with varying dependence.
+In a second set of experiments, both methods are evaluated for consistency tests with pairs of MNIST and CIFAR images.
+Because these do not have ground truth MI values, we exclude them from the table.
+In this work, a flow-based GM model is used for the first experiment, and a VAE-based GM model is used for the consistency experiment.
+MINDE (Franzese et al., 2023) used the BMI benchmark (Czyż et al., 2023) to evaluate their model.
+BMI contains 40 tasks.
+We count any unique combination of distribution and transform (including degrees of freedom and noise levels) to be a unique distribution type, excluding tasks that only differ by the dimensions or by correlation strength.
+Among the 40 tasks, there are 19 of these unique distributions, and 16 of these are non-normal.
+In their first experiment, they evaluate on these distributions with 100K samples, with MI ranging from 0.2 to 1.
+In a second experiment, they evaluate on 3x3 multivariate normal distributions, along with a half-cube and spiral transformation of them.
+Their experiment increases MI from 0 to 5.
+We count this experiment as having 3 unique distribution types, 2 of which are non-normal.
+The sample size is not directly reported.
+In a third experiment, they repeat the consistency tests from SMILE (Song and Ermon, 2019) on MNIST only, which we exclude because there is no ground truth MI.
+VCE Chen et al.
+(2025) first evaluate estimators on five synthetic distributions, under varying dependence.
 In this setting, MI ranges from 0 to 16, 10K samples are used, and dimensions range from 2 to 64.
-Then, the authors study increasing dimensionality on five distributions, with dimensions from 2 to 64 and MI values up to 14.
-10K samples were used, and six distributions across the two synthetic experiments were not gaussian.
-Next, authors consider correlated rectangles and gaussian plates.
+In a second experiment, the authors study increasing dimensionality on five distributions (three non-normal), with dimensions from 2 to 64 and MI values up to 14.
+10K samples were used.
+In a third experiment, authors consider correlated rectangles and gaussian plates.
 These images have dimension 16x16 but are processed by an autoencoder prior to prediction to produce 16 dimensional embeddings.
-The max MI for this setting is 7 and 10K samples were used.
-Finally, they consider language embeddings.
+The embedding portion is considered a pre-processing step, rather than part of the estimation method (as in LMI), and we therefore report the embedding dimension.
+This is consistent with other similar experiments, in which lower-dimensional data is scaled up to higher dimensions prior to evaluation, for which we correspondingly select the scaled dimensionality.
+The dimension reported is the dimension of data provided to the estimator.
+We consider this experiment to have 2 distributions, 1 of which is non normal, and dimension of 16.
+The maximum MI value is 7, and 10K samples used.
+In a fourth experiment, they consider language embeddings from Llama-3 13B and Bert.
+We consider this to be 2 non-normal distributions.
 Data is similarly processed by an autoencoder to be 16 dimensions.
-The maximum MI is 2.1, and 400 samples were used.
-MINE (mine) was first evaluated on 2-D and 20-D Gaussians with MI from 0 to 2 and 0 to 40, respectively.
+The underlying MI is estimated with resampling, and the maximum reported value is 2.1.
+4K samples were used.
+MINE (Belghazi et al., 2018) was first evaluated on 2-D and 20-D Gaussians with MI from 0 to 2 and 0 to 40, respectively.
 The number of samples used was not specified.
-In a second experiment, the authors applied a deterministic nonlinear transformation and demonstrated that their estimates are invariant to it.
-NWJ (nwj) introduced a variational bound for estimating the KL-divergence between two arbitrary distributions, and their experiments focused on evaluating the quality of this bound and estimation on various distributions.
+In a second experiment, the authors applied 3 deterministic nonlinear transformations to a normal distribution and demonstrated that their estimates are invariant to it.
+MINE then performs several downstream application experiments.
+NWJ (Nguyen et al., 2010) introduced a variational bound for estimating the KL-divergence between two arbitrary distributions, and their experiments focused on evaluating the quality of this bound and estimation on various distributions.
 The authors did not evaluate this approach on mutual information prediction directly, for which the distributions compared would be a joint distribution and the product of its marginals.
-InfoNCE (oord2018_cpc) was developed as a method for representation learning, with evaluation including representation learning tasks but without any evaluation of MI prediction itself.
+InfoNCE (Oord et al., 2018) was developed as a method for representation learning, with evaluation including representation learning tasks but without any evaluation of MI prediction itself.
+FMMI (Butakov et al., 2025) empirically evaluate their method on four distribution types (correlated normal, half-cube correlated normal, correlated uniform, smoothed uniform), with MI ranging from 0 to 80 nats.
+We convert nats (natural logarithm, base $e$) to bits (base 2), by dividing the nats value by $\log_{e}(2)$, and round to the nearest whole number (115).
+In this experiment, the training size was 100K samples, and the test set was 1K samples.
+The dimensions ranged from 0 to 160 dimensions.
+FMMI is also applied to predict O-information in fMRI data.
+InfoNet (Hu et al., 2024) perform their first experiment evaluating test-time efficiency on GMM distributions of various lengths, using sequence of 200, 500, 1000, 2000, and 5000 samples.
+Since no additional details about the estimation are included (maximum MI, dimension, number of samples), we exclude it from the table.
+In their second experiment (the first we report in the table), they benchmark MI estimation on 1D gaussian distributions with varying correlation.
+The MI ranges from 0 to 0.9, and 2000 samples are used.
+In the third experiment (the second we report in the table), the authors evaluate their method on 1D GMMs with multiple components (up to 20, which we consider as 20 distinct distribution types, 19 of which are non-gaussian), using 10 MI levels from 0 to 0.9.
+Authors use 2K samples.
+In their fourth experiment (the third we report in the table), the authors evaluate order accuracy on 1D GMMs with components from 1 to 10 (we count this as 10 distinct distributions, 9 of which are non-normal).
+The number of samples and maximum MI value are not reported.
+In their fifth experiment (the fourth we report in the table), they evaluate an application in high-dimensional independence testing, using slicing to scale their 1D-trained model.
+Their experiment tests three structures of correlation in gaussian distributions, and we report this as 3 distribution types, 0 of which are non-gaussian.
+They evaluate on 16 and 128 dimensions, and the number of samples varied from 50 to 950.
+In their sixth experiment, they evaluate their approach using MI to perform object segmentation in video data.
+Because this experiment is a downstream application without ground-truth MI values, we exclude it from the table.
+In this work, MIST, we use the BMI library to generate a range of 23 distribution types (16 in $\mathcal{M}_{\text{train}}$, 7 in $\mathcal{M}_{\text{test,OoMD}}$).
+Of these distributions, 19 are non-normal.
+The maximum dimension we use is 32, and we test MI values up to 40.
+The minimum sample size in our experiments is 10.
+Our downstream experiments in consistency tests and feature selection are excluded from this table, whose focus is evaluating MI estimation directly.
 Appendix B Details of Meta-Dataset Creation
-We construct a meta-dataset of synthetic distributions with known MI using the BMI library (bmi_benchmark), which generates complex yet tractable distributions via invertible transformations applied to base families with analytical MI.
+We construct a meta-dataset of synthetic distributions with known MI using the BMI library (Czyż et al., 2023), which generates complex yet tractable distributions via invertible transformations applied to base families with analytical MI.
 Each meta-sample consists of joint samples from a distribution paired with its ground-truth MI value, enabling supervised learning directly over distributions.
@@
 The true MI values for these distributions and their transformations are computed analytically or numerically, depending on tractability.
+Table 4: Meta-Dataset Description.
+Each distribution family is labeled as multi_<base_dist>-<structure>-<transform>, where <base_dist> denotes the base distribution type, <structure> specifies the parameterization structure, and <transform> indicates the nonlinear transformation applied to the data.
+For resource-constrained evaluation, a reduced test subset of 1k samples is used for both IMD and OoMD splits.
+The reduced subsets preserve the same distribution family composition as their full counterparts.
 Dataset Split Size (Extended / General) Distribution Families
@@
 multi_additive_noise-[base, halfcube]
-Table 4: Meta-Dataset Description.
-Each distribution family is labeled as multi_<base_dist>-<structure>-<transform>, where <base_dist> denotes the base distribution type, <structure> specifies the parameterization structure, and <transform> indicates the nonlinear transformation applied to the data.
-For resource-constrained evaluation, a reduced test subset of 1k samples is used for both IMD and OoMD splits.
-The reduced subsets preserve the same distribution family composition as their full counterparts.
 Appendix C Results
@@
 6, and (ii) predicting all four coefficients $\big(\hat{I}_{\mathrm{inv}},\hat{H}_{X,\mathrm{inv}},\hat{H}_{Y,\mathrm{inv}},\hat{H}_{XY,\mathrm{inv}}\big)$ in the multi-task setting.
+Figure 9: Distribution of true MI values within $\mathcal{M}_{\text{train}}$.
+Table 5: Mean MI estimates across prediction strategies on $\mathcal{M}_{\text{test}}$ ($\mathrm{dim}=9$), comparing normalization methods and coefficient configurations.
+“3-coef” uses the estimators $(\hat{H}_{X,\mathrm{inv}},\hat{H}_{Y,\mathrm{inv}},\hat{H}_{XY,\mathrm{inv}})$; the last two entries show predictions with $\hat{I}_{\mathrm{inv}}$, with and without the physics-informed loss $\mathcal{L}_{\mathrm{phys}}$.
+Values are (mean ± stddev) over three runs.
+Prediction Method $\mathcal{M}_{\text{test,IMD}}$ $\mathcal{M}_{\text{test,OoMD}}$
+Direct MI $\mathbf{5.4\mathrm{e}^{-2}}$ $\pm 2.0$ $\mathbf{1.2\mathrm{e}^{0}}$ $\pm 2.3$
+Log norm. ($\hat{I}_{\mathrm{log}}$) $8.5\mathrm{e}^{0}$ $\pm 2.1$ $1.4\mathrm{e}^{1}$ $\pm 2.4$
+Inverse norm. ($\hat{I}_{\mathrm{inv}}$) $8.7\mathrm{e}^{-2}$ $\pm 2.5$ $1.9\mathrm{e}^{0}$ $\pm 3.3$
+Entropy (3-coef) $2.5\mathrm{e}^{-1}$ $\pm 3.0$ $3.5\mathrm{e}^{0}$ $\pm 4.0$
+Entropy & $\hat{I}_{\mathrm{inv}}$ (with $\mathcal{L}_{\mathrm{phys}}$) $\mathbf{5.2\mathrm{e}^{-2}}$ $\pm 1.8$ $\mathbf{1.5\mathrm{e}^{0}}$ $\pm 2.0$
+Entropy & $\hat{I}_{\mathrm{inv}}$ (w/o $\mathcal{L}_{\mathrm{phys}}$) $7.0\mathrm{e}^{-2}$ $\pm 2.4$ $1.8\mathrm{e}^{0}$ $\pm 3.0$
 For the four-coefficient configuration, a physics-informed term was introduced to enforce consistency between two independent MI estimates.
@@
 The goal of this setup is to study whether predicting physically-related coefficients can help refine the model’s latent representations, incorporate additional structural knowledge, and improve prediction robustness.
-Figure 9: Distribution of true MI values within $\mathcal{M}_{\text{train}}$.
-Prediction Method $\mathcal{M}_{\text{test,IMD}}$ $\mathcal{M}_{\text{test,OoMD}}$
-Direct MI $\mathbf{5.4\mathrm{e}^{-2}}$ $\pm 2.0$ $\mathbf{1.2\mathrm{e}^{0}}$ $\pm 2.3$
-Log norm. ($\hat{I}_{\mathrm{log}}$) $8.5\mathrm{e}^{0}$ $\pm 2.1$ $1.4\mathrm{e}^{1}$ $\pm 2.4$
-Inverse norm. ($\hat{I}_{\mathrm{inv}}$) $8.7\mathrm{e}^{-2}$ $\pm 2.5$ $1.9\mathrm{e}^{0}$ $\pm 3.3$
-Entropy (3-coef) $2.5\mathrm{e}^{-1}$ $\pm 3.0$ $3.5\mathrm{e}^{0}$ $\pm 4.0$
-Entropy & $\hat{I}_{\mathrm{inv}}$ (with $\mathcal{L}_{\mathrm{phys}}$) $\mathbf{5.2\mathrm{e}^{-2}}$ $\pm 1.8$ $\mathbf{1.5\mathrm{e}^{0}}$ $\pm 2.0$
-Entropy & $\hat{I}_{\mathrm{inv}}$ (w/o $\mathcal{L}_{\mathrm{phys}}$) $7.0\mathrm{e}^{-2}$ $\pm 2.4$ $1.8\mathrm{e}^{0}$ $\pm 3.0$
-Figure 10: Mean MI estimates across prediction strategies on $\mathcal{M}_{\text{test}}$ ($\mathrm{dim}=9$), comparing normalization methods and coefficient configurations.
-“3-coef” uses the estimators $(\hat{H}_{X,\mathrm{inv}},\hat{H}_{Y,\mathrm{inv}},\hat{H}_{XY,\mathrm{inv}})$; the last two entries show predictions with $\hat{I}_{\mathrm{inv}}$, with and without the physics-informed loss $\mathcal{L}_{\mathrm{phys}}$.
-Values are (mean ± stddev) over three runs.
 Experiments on $\mathcal{M}_{\text{test}}$ with $\mathrm{dim}=9$ are summarized in Table 9.
@@
 C.2 Dimensionality Ablation
-Figures 11 through 13 contain additional bar plots for the experiment described in the second half of Section 4.6.
+Figures 10 and  11 contain additional bar plots for the experiment described in the second half of Section 4.7.
+We give in Figures 12 through 14 a high-level summary of the general performance of all various fixed- and variable-dimensionality models for ease of comparison.
 Figure 10: Average MSE obtained by variable- and fixed-dimensionality versions of MIST on subsets of $\mathcal{M}_{\text{test-extended,OoMD}}$ with fixed dimensions.
 […]
-[…]
-We give in Figures 14 through 16 a high-level summary of the general performance of all various fixed- and variable-dimensionality models for ease of comparison.
 Figure 12: Average MSE of the proposed models with variable input dimensionality versus dimension-specific variants.
@@
 […]
+Table 6: Mean mutual information (MI) estimates across data distributions.
+Color indicates whether distributions are from the in-meta-distribution (blue) or out-of-meta-distribution (red) subsets from $\mathcal{M}_{\text{test}}$.
+We also highlight in green values that are within one order of magnitude of the best and mark the best values in bold.
 CCA $7.7\mathrm{e}^{2}$ $1.1\mathrm{e}^{2}$ $2.0\mathrm{e}^{3}$ $1.8\mathrm{e}^{3}$ $1.3\mathrm{e}^{2}$ $9.1\mathrm{e}^{2}$ $1.2\mathrm{e}^{3}$ $1.5\mathrm{e}^{3}$ $8.8\mathrm{e}^{2}$ $1\mathrm{e}^{2} 1.1$$1\mathrm{e}^{3}$ $2.7\mathrm{e}^{2}$ $6.7\mathrm{e}^{2}$
@@
 Half-cube @ Additive Noise
-Table 5: Mean mutual information (MI) estimates across data distributions.
-Color indicates whether distributions are from the in-meta-distribution (blue) or out-of-meta-distribution (red) subsets from $\mathcal{M}_{\text{test}}$.
-We also highlight in green values that are within one order of magnitude of the best and mark the best values in bold.
 C.4 Inference Time
 We report in Table 7 the inference time of each of the baselines, averaged across the data from Figure 7.
+Table 7: Average inference time per sample measured on $\mathcal{M}_{\text{test-extended}}$.
 Method Time / sample (s)
@@
 LMI $0.76$
+InfoNet $1.64$
 InfoNCE $2.01$
@@
 MINE $2.75$
+VCE $3.93$
+FMMI $15.71$
 MINDE $122.7$
-Table 6: Average inference time per sample measured on $\mathcal{M}_{\text{test-extended}}$.
 C.5 Scaling Behavior
-We report in Table 7 the performance of the main model (5.5M) used in our experiments compared to scaled-down models.
+We report in Table 8 the performance of the main model (5.5M) used in our experiments compared to scaled-down models.We give in Figure 15 a fine-grained heatmap of the MSE performance obtained by MIST and MISTQR for various sample size groups and dimensions.
+Table 8: Average performance metrics for models of different sizes.
 Model Params IMD OoMD
@@
 […]
-Table 7: Average performance metrics for models of different sizes.
-We give in Figure 17 a fine-grained heatmap of the MSE performance obtained by MIST and MISTQR for various sample size groups and dimensions.
 Figure 15: Detailed visualization of the average MSE values for our two proposed models across different sample-size groups and all evaluated dimensionalities.
@@
 When moving to OoMD distributions, however, those failure rates increase to 15.8% for lower bounds and 27.5% for upper ones, showing that this task remains difficult for our model.
+[…]
 IMD (1080 points) OoMD (1260 points)
@@
 Upper bound failures 46 (4.3%) 346 (27.5%)
+D.2 Self-Consistency of the Learned Quantity
 […]
-D.2 Self-Consistency of the Learnt Quantity
-[…]
-  • The independence test checks that if $X\mathrel{\perp\!\!\!\perp}Y$, then $I(X;Y)=0$.
+  • The independence test checks that if $X\perp\!\!\!\!\perp Y$, then $I(X;Y)=0$.
   • The data processing inequality checks that if $X,Y$ and $Z$ form a Markov chain $X\to Y\to Z$, then $I(X;Y)\geq I(X;Z)$ (further processing of the data cannot add new information).
-  • The additivity over independent samples test checks that if $(X_{1},X_{2})\mathrel{\perp\!\!\!\perp}(Y_{1},Y_{2})$, then $I(X_{1},Y_{1};X_{1},Y_{2})=I(X_{1};X_{2})+I(Y_{1},Y_{2})$
+  • The additivity over independent samples test checks that if $(X_{1},X_{2})\perp\!\!\!\!\perp(Y_{1},Y_{2})$, then $I(X_{1},Y_{1};X_{1},Y_{2})=I(X_{1};X_{2})+I(Y_{1},Y_{2})$
 (a) MSE Model
@@
 Figure 17: Data Processing Test: AA is an MNIST image, and $B_{r}$ contains the top $r$ rows of image AA.
-This test compares $I([A,A],[B_{r},B_{{}^{\prime}r-3^{\prime}}])$ to $I([A,B_{r}])$, where AA is an original MNIST image, $B_{r}$ is the first $r$ rows of $A$, and so $B_{{}^{\prime}r-3^{\prime}}$ is a further masked version of image AA.
+This test compares $I([A,A],[B_{r},B_{{}^{\prime}r-3^{\prime}}])$ to $I([A,B_{r}])$, where $A$ is an original MNIST image, $B_{r}$ is the first $r$ rows of $A$, and so $B_{{}^{\prime}r-3^{\prime}}$ is a further masked version of image AA.
 The test produces a ratio which should be near to 1 for all values of $r$, which amounts to checking that adding a further masked version of your image $B_{r}$ does not change the predicted MI.
@@
 Error bars of $1\sigma$ are reported with 100 bootstrapped samples.
+Appendix E Feature Selection
+In this section, we explore using our estimators for a simple feature selection problem in which feature weights are learned by maximizing the mutual information with an outcome Y.
+We define $X\in\mathbb{R}^{n\times d}$ as a random normal matrix with $d$ features and $n$ samples.
+Next, we define a set of weights for these features, each bounded between 0 and 1, $w\in[0,1]^{d\times 1}$.
+The label Y is given by $Y_{\text{true}}=Xw^{*}+\epsilon$, where $Y\in\mathbb{R}^{n\times 1}$ is the matrix multiplication of $X$ and pre-set masking weights $w^{*}$ plus added normal noise $\epsilon$.
+For a candidate set of weights, the predicted label is $Y^{\prime}=Xw^{\prime}\in\mathbb{R}^{n\times 1}$, produced by the candidate weights $w^{\prime}$.
+The best weights are chosen as $w_{\text{best}}=\operatorname*{arg\,max}_{w^{\prime}}\text{MIST}(Y_{\text{true}},Y^{\prime})$, which maximizes the mutual information with the true label.
+The goal of this task is to recover the true weights ($w^{*}$), which correctly assign a value of 1 for the two informative features and 0 otherwise.
+We set the mean and variance of $X$ to 0 and 1.0 respectively, and the normal noise applied to $Y_{true}$ has 0 mean and variance of either 0.3, 0.5, or 1.0.
+We generate 32 features and randomly select 2 features to be informative.
+We use MIST and MINE estimators, and perform gradient ascent with the Adam optimizer and learning rate of 0.01 for 500 steps.
+For MINE, we consider two settings.
+In the first, we continually train one MINE estimator, allowing 15 steps of updates with each prediction.
+In the second setting, we train a new MINE estimator for 200 steps for each prediction.
+In both cases, we follow the parameters in MINE’s repository, using Adam for the optimizer and a learning rate of 0.001.
+We used the biased version loss and set the critic hidden dimension to 64.
+Weights are initialized with a small constant and bounded between 0 and 1.
+In Figure 19, we plot the ROC curves of the learned features, which includes 3 noise values (0.3, 0.7, and 1.0) and 20 random seeds each.
+We find that even with a very limited number of samples, MIST estimates provide an accurate and stable learning signal.
+Training MINE to convergence at each prediction achieves better performance than continually MINE training, but remains significantly less effective than using MIST estimates.
+In Figure 20, we plot the top-k accuracy according to the number of samples provided.
+Accuracy is reported as the proportion of the two highest weighted features that are truly informative.
+While the number of informative features is rarely known ahead of time, we provide this view of the results to better compare the method performance at each sample size.
+We find that MIST-guided selection is more accurate and lower variance on average, achieving perfect accuracy at 100 samples.
+In Figure 21, we show the feature weights through training of two randomly chosen seeds with noise variance set to 1.0.
+Figure 19: ROC curves of feature selection learned with MINE and MIST estimates.
+Results are computed with 3 noise levels and 20 random seeds, with 1 standard deviation shaded.
+We find that MIST-guided feature selecting is more accurate and lower variance across settings.
+Figure 20: Comparison of the accuracy of features selected via gradient ascent on MINE predictions (either trained continually and trained to convergence) and MIST predictions.
+Accuracy is calculated as the proportion of the highest two weights which are the informative features.
+In other words, how often are top 2 weights assigned to the two correlated features?
+We find that MIST-guided feature selection is more accurate and lower variance.
+MINE (Continual)
+MINE (Convergence)
+MIST
+Figure 21: Comparison of feature selection guided by MINE with continual training (left), MINE with convergence training (middle), and MIST (right).
+Plotted are the feature weights throughout training on two randomly selected seeds of data (rows).
+We find MIST-guided training to be very effective at differentiating correlated features.